### v2

# M3RS: Multi-robot, Multi-objective, and Multi-mode Routing and Scheduling

Abstract
The quality of task execution can significantly impact a multi-robot mission. While higher quality is desirable, it may not always be feasible due to mission constraints. Existing multi-robot task allocation literature generally overlooks quality of service as a decision variable. Addressing this gap, we introduce the multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem, designed for time-bound, multi-robot, multi-objective missions. In M3RS, each task offers multiple execution modes, each with different resource requirements, execution time, and quality. M3RS optimizes task sequences and execution modes for each agent. The need for M3RS comes from multi-robot applications in which a trade-off between multiple criteria can be achieved by varying the task level quality of service through task execution modes. Such ability is particularly useful for service robot applications. We use M3RS for the application of multi-robot disinfection in healthcare environments and other public locations. The objectives considered for disinfection application are disinfection quality and number of tasks completed. A mixed-integer linear programming (MIP) model is proposed for M3RS. Further, a clustering-based column generation (CCG) algorithm is proposed to handle larger problem instances. Through synthetic, simulated, and hardware case studies, we demonstrate the advantages of M3RS, showing it provides flexibility and strong performance across multiple metrics. Our CCG algorithm generates solutions 2.5$\times$ faster than a baseline MIP optimizer, maintaining competitive performance. The videos for the experiments are available on the project website [^1] https://sites.google.com/view/g-robot/m3rs/.
Index Terms:
multi-robot system, autonomous systems, task allocation, scheduling, service robots, disinfection robots.
I Introduction and Related Works
Maintaining hygiene in healthcare environments is crucial to prevent catastrophic issues like hospital-acquired infections. Disinfection robots are useful technologies that provide an additional layer of protection by automating the disinfection of target surfaces. A team of these robots can improve both spatial and temporal efficiency in maintaining hygiene standards in hospitals [1]. To coordinate a fleet of robots for a disinfection mission, two key decisions must be made: i) which task will be assigned to which robot, and ii) what level of disinfection quality should be applied to each task. While higher disinfection quality is desirable, it also requires more time and resources, which may not always be feasible. The assignment of tasks to robots has been studied in the literature [2], but there is a lack of research addressing the decision regarding the quality of service of a task. In disinfection applications, considering the quality of tasks is important as it impacts overall hygiene standards. Additionally, some surfaces, such as those in patient rooms and ICUs, require stringent quality standards. In this work, we tackle the problem of deciding which tasks to implement and at what mode/quality.
For a team of robots, Multi-robot Task Allocation (MRTA) algorithms [3, 2] allocate tasks to various robots working on a common mission while optimizing a joint mission objective. A key metric to evaluate these allocations is the quality of service (QoS) for the mission. The definition of QoS depends on the mission requirements and target application, such as the quality of communication between the robot and the base station [4, 5], the tardiness of addressing tasks [6], the average waiting time for a customer [7], or productivity and fatigue [8]. Often, there is a trade-off between QoS and other mission objectives, which could be conflicting. In such cases, multiple solutions exist that prioritize these objectives [4, 5, 8].
Figure 1: M3RS is suitable for routing and scheduling in time-bound scenarios where multiple criteria of interest depend on the mode in which the agents execute the tasks. This visualization depicts two cases. Here, mode 1 stands for the highest quality and long task execution times, whereas mode 3 is for the lowest quality and fastest task execution times. The left side of the image shows a case where agents choose to maximize the number of tasks completed by performing tasks in lower-quality modes. On the right, agents prioritize quality over quantity. As a result, tasks 2 and 5 are missed. An example of such a scenario is a multi-robot system performing services such as disinfection and cleaning.
MRTA problems are $\mathcal{NP}$-hard, making larger instances challenging to solve. With extra variables, concerning the quality of the service, the problem becomes more complex. Techniques such as heuristics and metaheuristics are faster but often sacrifice optimality [9]. Learning-based methods require expensive training and massive datasets, and may still struggle to generalize [10]. Column generation is a decomposition approach that refines the solution iteratively by solving smaller sub-problems. This allows faster convergence to higher-quality solutions as it directly exploits the problem’s structure and efficiently handles large-scale instances. We propose a clustering-based column generation algorithm, where clustering further helps in reducing the search space to find good quality solutions.
To the best of our knowledge, this is the first work to address task-level service quality using multiple execution modes for quality-dependent multi-robot missions. We introduce a multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem. The contributions of this work are:
  1. 1.
    We present M3RS, a novel problem to alter task-level QoS to achieve trade-offs in different mission objectives. The proposed problem is well-suited for applications where tasks are spread across a physical space, each task has service requirements and resource requirements, and the overall mission has multiple criteria of interest. Each robot can address the task in multiple execution modes, where each mode corresponds to different task quality, resource consumption, and service times. A visual depiction of the application of M3RS is given in Fig. 1.
  2. 2.
    We present the usage of M3RS for a multi-robot disinfection application. A thorough study of different mission scenarios is conducted on a synthetic dataset. Further, we test M3RS in simulated and hardware case studies.
  3. 3.
    We use a mixed-integer linear programming formulation for the proposed problem. To handle the large decision space, we propose clustering-based column generation (CCG) for solving M3RS. CCG generates solutions 2.5$\times$ faster than the baseline while maintaining competitive performance.
This paper is organized as follows: The general formulation and optimization of M3RS is given in Section III. The results of the synthetic, simulated, and real-world experiments are presented in Section IV. Finally, Section V includes future works and conclusions.
II Literature Review: QoS in Task Assignemnt
QoS has also received attention in routing and production scheduling literature. Exposito et al. [11] considered the timeliness of vehicles as a QoS indicator for vehicle routing. Similarly, QoS has been studied for home healthcare applications, where it involves routing nurses to different patients for healthcare-related services, depending on the patient receiving care from a nurse with the desired experience [12, 13]. In supply chain and production planning, QoS has been considered as the existence of minimum stock [14]. Most of these existing works modify the task allocation to achieve different trade-offs between QoS and other objectives. However, there are no works that focus on the deliberate degradation or improvement of QoS offered to an individual task by a robot as a means to achieve an overall trade-off for various mission objectives.
Task execution modes are different operational settings for servicing tasks. These modes depend on various constraints in the environment or aim to achieve desired performance on key metrics like quality and resource consumption. Lahijanian et al. [15] explored using an energy-saving mode where resource-intensive autonomy features are disabled to conserve energy in resource-constrained environments. The usage of task execution modes in MRTA has been limited. Task execution modes were used to decide if humans and robots should collaborate to maximize time efficiency [16, 17]. In project scheduling, task execution modes are used to minimize project makespan while adhering to resource constraints [18, 19, 20]. In vehicle routing, mode selection involves choosing delivery locations with specific constraints, impacting customer delivery [21]. The hybrid electric vehicle routing problem in [22] used mode selection for switching driving modes, affecting emissions and transportation costs. Using task execution modes to select QoS provided to the task is still unexplored. This decision affects the quality of individual tasks and the overall mission. For example, in a disinfection mission, cleaning a surface with the highest available dosage ensures the highest quality level. However, higher dosages require longer service times, which could potentially result in skipping other tasks.
III Multi-Robot Multi-Objective Multi-Mode Routing and Scheduling
In this section, first, the general optimization model for M3RS problem is presented. Then we present a column generation (CG) scheme to solve the optimization model with faster computation times. Our novel contribution is the use of clustering heuristics, which helps solve the problem more quickly and find solutions with better objective values. Finally, the formulation is applied to the multi-robot disinfection application.
Table I: Notation for M3RS.
Notation Description
$\mathcal{A}$ Set of agents.
$\mathcal{T}$ Set of tasks.
$\mathcal{M}_{i}$ Set of modes for task $i$, where $\mathcal{M}_{i}=\{1,2...M\}$.
$R$ Number of objective criteria.
$f_{r}$ $r^{th}$ objective criterion.
$p_{i,m}$ The quality indicator of doing task $i$ in mode $m$.
$p^{max}$ The highest possible quality of disinfection
$\sigma_{i,m}$ Time to complete task $i$ in mode $m$.
$q_{i\scriptsize{\xrightarrow{}}j}$ Energy consumed for travelling from task $i$ to task $j$.
$q_{idle}^{k}$ Energy consumption during idling by agent $k$.
$q_{i,m}$ Energy consumption of task $i$ in mode $m$.
$\text{q}_{\text{travel}}$ Avg. current consumption in traveling at a constant speed.
$\text{q}_{\text{tc}}$ Avg. current consumption while completing tasks.
$\text{q}_{\text{idling}}$ Avg. current consumption while idling.
$[a_{i},b_{i}]$ Start and end time for task $i$.
$Q$ Total capacity of an agent.
$T_{H}$ Time available to complete all tasks.
$t_{i,j}$ Travel time from task $i$ to task $j$.
$\lambda_{r}$ The scalar preference for $r^{th}$ objective function.
$s^{k}_{i}$ A decision variable for arrival time of agent $k$ to task $i$.
$x^{k}_{i{\xrightarrow{}}j}$ A binary decision variable for if agent $k$ goes from task $i$ to task $j$.
$y^{k}_{i,m}$ A binary decision variable for task $i$ done in mode $m$ by agent $k$.
$\eta$ The load factor hat is used to simulate different disinfection missions.
III-A Optimization Model
A solution for the M3RS problem determines routes, schedules, and task execution modes for a set of agents $\mathcal{A}$, and a set of tasks $\mathcal{T}$ distributed across physical space. Refer to Table I for notation. All the agents are identical, and only one agent can service each task. We use $i$ and $j$ to denote different tasks in $\mathcal{T}$ and $k$ to denote $k^{th}$ agent in the rest of the notation. The time window for a task $i$ specifies a starting time, $a_{i}$, and finishing time, $b_{i}$, within which the task must be serviced. The entire multi-robot operation must be done in a fixed mission time of $T_{H}$ hours. Further, each task $i$ has an associated mode set $\mathcal{M}_{i}=\{1,2...M\}$, representing available task service modes. Each task can have distinct modes of service available, so each task has a distinct mode set. An agent services the task in one of the modes available in the associated mode set. The selected mode $m\in\mathcal{M}_{i}$ for task $i$ has a corresponding service time ($\sigma_{i,m}$) and energy requirement ($q_{i,m}$) from the agent. Performing a task in a mode with higher service time and resource consumption would lead to lower availability of agents for other tasks. The overall scheduling and allocation mission is optimized for multiple objectives.
The energy capacity for each agent is given by $Q$ Ah (Ampere-hours). The energy consumption for each agent is divided into performing tasks, traveling, and idling [23]. A fixed-rate battery model is used to calculate the energy consumption, where energy consumed is proportional to idling time, task execution time, and traveling time respectively [23]. Different aspects of energy expenditure can be calculated using:
$\begin{split}q_{i\scriptsize{\xrightarrow{}}j}=\text{q}_{\text{travel}}\cdot t_{i,j},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. traveling energy)}\\ q_{i,m}=\text{q}_{\text{tc}}\cdot\sigma_{i,m},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. servicing energy)}\\ q_{idle}^{k}=\text{q}_{\text{idle}}\cdot(M_{k}-W_{k}).\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. idling energy)}\end{split}$ (1)
Here, $\text{q}_{\text{travel}}$ is the current consumption for an agent traveling at a constant speed, $\text{q}_{\text{tc}}$ is the current consumption while performing a task, and $\text{q}_{\text{idle}}$ is the current consumption for idling. $M_{k}$ is the total time spent by agent $k$ on the mission, and $W_{k}$ is the total time spent by agent $k$ traveling and executing tasks. The net energy consumed by agent $k$ in a mission is given by the LHS of Eq. (9). The relevant notation for the formulation is summarised in Table I.
The formulation for M3RS is expressed as a mixed integer with the following decision variables:
i) the binary decision variable $x^{k}_{i{\xrightarrow{}}j}\in\{0,1\}$ for assignment and transition of agent $k$ from task $i$ to $j$;
ii) the binary decision variable $y^{k}_{i,m}\in\{0,1\}$ for allocation of mode $m$ for task $i$ when executed by agent $k$; and
iii) the real continuous variable $s^{k}_{i}\in\mathbb{R}$ denotes arrival time of agent $k$ to task $i$.
The optimization problem is expressed as
$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m},s^{k}_{i}}{\max}\,\sum_{r=0}^{R}\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m},s^{k}_{i}),$ (2)
subject to the constraints in Eq. (3)-(16).
Eq. (2) is the scalar objective function that maximizes the convex combination of $R$ objective criteria. Here, $f_{r}$ is $r^{th}$ objective criteria for $r=1,2,...R$. The user-specified parameter $\lambda_{r}$ represents trade-off preference for $r^{th}$ objective criterion. Each objective criterion can depend on task allocation, selected mode, and arrival times.
The constraints of the optimization are categorized into five sets as follows: 1) task assignment constraint, 2) feasibility constraints, 3) energy consumption constraint, 4) time constraints, and 5) domain and preference constraints. These constraints are described next.
III-A1 Task Assignment Constraint
Given by Eq. (3), the task assignment constraint ensures that each task is done at most once:
[…]
Here, $i=0$ is excluded ($\mathcal{T}\backslash 0$) as it corresponds to the home depot. This is done to maintain feasibility as all agents must leave the home depot. There might be missions where some tasks must be completed. For such cases, the inequality can be replaced by equality for tasks that must be completed.
III-A2 Feasibility Constraints
Given by Eq. (4)-(8), the feasibility constraints ensure the solutions are logical.
[…]
The constraint in Eq. (4) eliminates infeasible solutions where an agent transitions between the same task. The constraints in Eq. (5)-(6) ensure that all agents leave the home depot and return to the home depot, respectively. The constraint in Eq. (7) ensures that the agent must exit the tasks they visit. The constraint in Eq. (8) enforces that each task $i$ can be performed in only one of the modes in $\mathcal{M}_{i}$. Thus, for any task $i$ selected by agent $k$, only one execution mode is used.
III-A3 Energy Consumption Constraint
As described by Eq. (9), the energy consumption constraint enforces that the energy expenditure of an agent in the form of traveling, doing tasks, and idling does not exceed the agent’s capacity:
$\text{q}_{\text{idle}}^{k}+\sum_{i\in\mathcal{T}}(\sum_{m\in\mathcal{M}_{i}}q_{i,m}\cdot y^{k}_{i,m}+\sum_{j\in\mathcal{T}}q_{i\scriptsize{\xrightarrow{}}j}\cdot x^{k}_{i{\xrightarrow{}}j})\,\leq\,Q\,\,\,\forall k\in\mathcal{A}.$ (9)
III-A4 Time Constraints
Given by Eq. (10)-(12), the time constraints ensure the timing of the tasks are consistent:
$s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}+t_{i,j}-s^{k}_{j}\leq 0\,\,\forall i\in\mathcal{T}$ (10)
$\forall j\in\mathcal{T}\,\,\forall k\in\mathcal{A},$
$a_{i}\leq s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}<b_{i}\,\,\,\forall i\in\mathcal{T},\,\,\forall k\in\mathcal{A},$ (11)
$0\leq s^{k}_{0}<T_{H}\,\,\forall k\in\mathcal{A}.$ (12)
The constraint in Eq. (10) ensures that the arrival time for a subsequent task is always later than the previous task. This constraint also enables sub-tour elimination [24], where sub-tours are infeasible solutions that represent disjoint sequences of completing tasks for the same agent. The constraint in Eq. (11) ensures that the agent arrives and completes the task in the selected mode within the task’s prescribed time window. Finally, the constraint in Eq. (12) ensures that all agents return to the depot within the specified time horizon $T_{H}$.
III-A5 Domain and User Preference Constraints
Given by Eq. (13)-(16), the preference constraints ensure the variable and parameters follow the definitions:
$x^{k}_{i{\xrightarrow{}}j}\in\{0,1\},$ (13)
$y^{k}_{i,m}\in\{0,1\}.$ (14)
$0<\lambda_{r}\leq 1,$ (15)
$\sum_{r=0}^{R}\lambda_{r}=1.$ (16)
The constraints in Eq. (13)-(14) enforce binary values for $x^{k}_{i{\xrightarrow{}}j}$ and $y^{k}_{i,m}$, while Eq. (15)-(16) ensure convexity according to given user preferences.
III-B A Solution for M3RS: Clustering-based Column Generation
M3RS extends traditional routing and scheduling problems by adding the decision of mode selection, which impacts task quality. Routing and scheduling problems are known to be $\mathcal{NP}-$hard [25]. Consequently, larger problem instances with many tasks and agents lead to exponential growth in the solution space and significantly burden computational resources [26]. We use CG [27, 28], to address the computational challenges associated with the large decision variables in the proposed formulation. The strength of CG lies in its ability to break down a complex problem into manageable sub-problems while iteratively refining the solution.
The CG algorithm decomposes the original optimization problem into a master problem and a sub-problem. The master problem identifies optimal columns (or paths/solutions) from a pool of feasible candidate solutions, $\Omega$. Since $\Omega$ is very large, typically, a restricted master problem is solved on a subset $\Omega^{\prime}\subset\Omega$. The sub-problem uses the dual solution from the master problem to find new favorable columns, which are added to $\Omega^{\prime}$ to improve the objective value of the master problem. The master problem and sub-problem are solved iteratively until there are no new columns with positive reduced cost. We refer to our approach as CCG (clustering-based CG) due to the use of clustering-based heuristics for solving the sub-problem. The rest of the section presents the building blocks of our CCG algorithm.
III-B1 Restricted Master Problem
This problem aims to identify the best set of solutions from the existing solution pool, resembling a set covering problem. Therefore, to solve the restricted master problem, we adopt a set covering problem formulation [27]. A solution for the set covering problem selects a subset of sets (in our case, task sequence and modes for each agent) from a collection (the solution pool $\Omega^{\prime}$) such that most of the elements (tasks) are covered. This must be done while adhering to constraints, such as each task being assigned to only one agent, and maximizing the given objective. The set covering formulation-based restricted master problem finds optimal solutions from a subset of feasible candidate solutions $\Omega^{\prime}$. Each candidate in $\Omega^{\prime}$ is a feasible solution that specifies a sequence of addressing different tasks in the associated modes and arrival times. The pool $\Omega^{\prime}$ can be initialized by some trivial solutions and can be updated iteratively with better solutions obtained by solving the sub-problem. The set covering formulation for our restricted master problem is given by:
$\underset{\theta^{k}}{\max}\,\sum_{k\in\Omega^{\prime}}c_{k}\cdot\theta^{k}$ (17)
subject to:
$\begin{split}\sum_{k\in\Omega^{\prime}}a^{k}_{i}\cdot\theta^{k}\leq 1\,\,\forall i\in\mathcal{T}/\{0\},\\ \end{split}$ (18)
$\sum_{k\in\Omega^{\prime}}b^{k}_{i,m}\cdot\theta^{k}\leq 1\,\,\forall i\in\mathcal{T}/\{0\},\forall m\in\mathcal{M}_{i}$ (19)
$\sum_{k\in\Omega^{\prime}}\theta^{k}\leq|\mathcal{A}|,$ (20)
$\theta^{k}\in\{0,1\}$ (21)
Here, $\theta^{k}$ is a binary solution selection variable, where $\theta^{k}=1$ if solution $k\in\Omega^{\prime}$ is selected, and $\theta^{k}=0$ otherwise. The objective function in Eq. (17) selects the candidates from $\Omega^{\prime}$ that maximizes the original objective function Eq. (23). The constraint in Eq. (18) ensures that each task in $\mathcal{T}$ is assigned at most once. Here, $a^{k}_{i}$ is 1 if $i^{th}$ task is covered in $k^{th}$ solution. The constraint in Eq. (19) ensures that mode $m$ for task $i$ is used atmost once with $b^{k}_{i,m}$ being 1 if $i^{th}$ task in $k^{th}$ solution is done in mode $m$. The main purpose of Eq. (19) is to identify the modes to explore via dual values. Finally, Eq. (20) ensures the number of candidates selected from $\Omega^{\prime}$ does not exceed the number of available agents. We apply a linear relaxation to the integer program defined in Eq. (17)-(21) by relaxing the constraint in Eq. (21) and allowing $\theta^{k}\in[0,1]$. This relaxation enables faster computation, as algorithms like the simplex method [29] can efficiently solve large-scale linear programs in polynomial time. This is especially useful with a large number of columns in $\Omega^{\prime}$.
III-B2 Clustering-based Sub-problem
In CG, the sub-problem is solved iteratively to populate $\Omega^{\prime}$. Solving the sub-problem efficiently with high-quality candidate solutions enhances the CG’s exploration of the solution space and accelerates convergence to better solutions. However, as the number of tasks and agents increases, the expanded solution space slows down the search for promising candidates. To address this, we introduce a clustering method inspired by cluster-first, route-second heuristics commonly used in routing and scheduling [30]. Our sub-problem solves the MIP problem in Eq. (22) for a single agent at a time, on a restricted set of tasks $\mathcal{T}^{\prime}$. All tasks in $\mathcal{T}$ are clustered based on time windows. We use a K-Means clustering [31] approach to create $n_{c}$ clusters, with $n_{c}$ being a user-defined parameter. Then $\mathcal{T^{\prime}}$ is generated by uniformly sampling each cluster. These modifications allow CCG to converge to higher-quality solutions faster than the baseline, as demonstrated in Section IV.
The sub-problem creates a solution for a single agent setup with a restricted task pool $\mathcal{T}^{\prime}$. The formulation for the sub-problem is:
$\begin{split}\underset{x_{i\scriptsize{\xrightarrow{}}j},y_{i,m}}{\max}\,\sum_{r=0}^{R}\sum_{i\in\mathcal{T}^{\prime}}\sum_{j\in\mathcal{T}^{\prime}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x_{i\scriptsize{\xrightarrow{}}j},y_{i,m},s_{i})-\\ \sum_{i\in\mathcal{T}^{\prime}/\{0\}}(\gamma_{i}\cdot\sum_{j\in\mathcal{T}^{\prime}}x_{i\scriptsize{\xrightarrow{}}j})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}\sum_{m\in\mathcal{M}_{i}}(\beta_{i}\cdot y_{i,m}),\end{split}$ (22)
subject to constraints Eq. (3)-(16) and the number of agents is restricted to one. The objective function of the sub-problem, Eq. (22), is a modified variant of Eq. (2), tailored for a single-agent setup and augmented with dual variables from the restricted master problem. Here, $\gamma_{i}$ and $\beta_{i,m}$ are non-negative dual of the restricted master problem corresponding to constraints Eq. (18) and Eq. (19) respectively. These dual variables are used to identify promising new candidate solutions. It should be noted that sub-problem The solution of the sub-problem is added to $\Omega^{\prime}$.
Algorithm 1 The clustering-based column generation algorithm
[…]
2:  Output: Selected solution for each agent from $\Omega^{\prime}$
3:  $\mathcal{T}_{c}^{init}\leftarrow$ Divide tasks in $\mathcal{T}$ to $n_{c}$ clusters based on time-windows using K-Means clustering.
4:  $\Omega^{\prime}\leftarrow$ Initialize the candidate pool
5:  while True do
6:     $\theta^{k},\gamma_{i},\beta_{i,m}\leftarrow$ Solve Eq. (17)-(20) subject to $\theta^{k}\in[0,1]$
7:     $\mathcal{T}_{v}\leftarrow$ Initialize empty list to store visited customers.
8:     for $i=1$ to $|\mathcal{A}|$ do
9:        if $i=1$ then
10:           $\mathcal{T}_{c}\leftarrow\mathcal{T}_{c}^{init}$
11:        else
12:           $\mathcal{T}_{c}\leftarrow$ Create $n_{c}$ clusters of remaining tasks ($\mathcal{T}/\mathcal{T}_{v}$) using K-Means clustering.
13:        end if
14:        $\mathcal{T^{\prime}}\leftarrow$ Uniformly sample tasks from $\mathcal{T}_{c}$
15:        $\mathcal{S}\leftarrow$ Solve the sub-problem using Eq. (22)
16:        if $\mathcal{S}$ is not empty then
17:           $\Omega^{\prime}\leftarrow$ Update solution pool with all feasible solutions from $\mathcal{S}$
18:           Update $\mathcal{T}_{v}$ with tasks addressed in the best solution of $\mathcal{S}$
19:        end if
20:     end for
21:     if time-limit reached or max-iterations completed then
22:        Terminate
23:     end if
24:  end while
25:  if Terminate then
26:     Solve Eq (17)-(20) subject to $\theta^{k}=\{0,1\}$
27:  end if
28:  return Selected solution for each agent from $\Omega^{\prime}$
III-B3 Complete Algorithm
The CCG algorithm is detailed in Alg. 1. In line 1, the algorithm receives problem parameters, user preferences, and key algorithm settings, including the number of clusters ($n_{c}$), the size of the restricted task set ($|\mathcal{T^{\prime}}|$), maximum iterations, and the time limit. In line 2, initial task clusters $\mathcal{T}_{c}^{init}$ are formed by dividing $\mathcal{T}$ into $n_{c}$ clusters based on time windows using K-Means. In line 3, the candidate solution pool $\Omega^{\prime}$ is initialized, with initial solutions generated by randomly selecting tasks and modes while ensuring constraints in Eq. (3)-(16) are satisfied.
From line 5, the restricted master problem and sub-problem are solved iteratively to generate new solutions. The restricted master problem is updated in each iteration based on $\Omega^{\prime}$ (line 6), and the sub-problem’s objective is guided by the duals obtained from the restricted master. A list $\mathcal{T}_{v}$ is initialized to track visited tasks (line 7). The sub-problem is solved iteratively for each agent in $\mathcal{A}$ (lines 8-20).
In the first iteration, task clusters $\mathcal{T}_{c}$ are initialized to $\mathcal{T}_{c}^{init}$ (line 10). For subsequent iterations, clusters are recreated from the remaining tasks (line 12). The restricted task set $\mathcal{T^{\prime}}$ is formed by uniformly sampling clusters from $\mathcal{T}_{c}$ (line 14), after which the sub-problem is solved (line 15). Feasible solutions with positive reduced costs are added to $\Omega^{\prime}$ (line 17), and $\mathcal{T}_{v}$ is updated with tasks corresponding to the best solutions in $\mathcal{S}$ (line 18). This process continues until the stopping criteria—maximum iterations, time limit, or lack of solutions with positive reduced costs—are met (lines LABEL:lst:ln21-23). Upon termination, the restricted master problem is solved to optimality with integrality constraints to obtain the final solution (lines 25-28). A time limit is applied to both the restricted master and sub-problems to maintain time efficiency throughout the process.
III-C M3RS for Multi-Robot Disinfection
We consider a fleet of surface disinfection robots deployed in hospital-like environments to provide around-the-clock disinfection. Specifically, we use G-robots [32], which are far-UV-based mobile manipulators designed for surface disinfection. A disinfection mission is a collection of disinfection tasks spread in an environment such as a hospital. Each mission has a specified time horizon within which it should be completed. Further, the tasks must be done in their specified time window. The tasks can have different disinfection dose requirements. This dosage dictates disinfection quality, energy consumption, and the task service time. Each task can be disinfected in a particular set of modes, where the modes have different disinfection dosage options for that task.
The disinfection dosage for UV-based surface disinfection is product of UV exposure duration and irradiance of the UV light source [1]. A higher disinfection dosage would require a longer exposure duration. Typically, the disinfection dosage is expressed in terms of percentage of population reduction. For example, a dosage of $D_{90}$ implies $90\%$ reduction of pathogenic population. A higher reduction percentage requires stronger disinfection dose which in turn leads to longer exposure times. Longer exposure times further implies higher energy consumption. We consider four disinfection modes that are $D_{99.9999}$, $D_{99.99}$, $D_{99}$, and $D_{90}$. Here, $D_{99.9999}$ is the strongest disinfection dose with the highest disinfection time and energy consumption as compared to $D_{90}$ [1]. The set of modes can be determined by the acceptable disinfection levels for a surface. So, each task has an associated set of acceptable disinfect doses or modes. The assignment of robots to different tasks and the corresponding modes will be determined by solving M3RS.
For disinfection applications, we have two objectives: i) the number of tasks completed and ii) overall disinfection quality. Higher disinfection quality requires more time and resources, limiting availability for other tasks. For example, critical locations like ICUs in hospitals require higher dosages, so higher quality is preferable. In contrast, areas like waiting rooms or offices do not need the highest quality, and covering the maximum number of surfaces might be preferable. Our formulation creates schedules by using user-specified trade-off requirements. The formulation of M3RS for disinfection application is:
$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m}}{\max}\,\frac{1}{|\mathcal{T}|}\cdot\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}(\lambda)\cdot x^{k}_{i{\xrightarrow{}}j}+(1-\lambda)\cdot\frac{p_{i,m}}{p^{max}}\cdot y^{k}_{i,m,}$ (23)
subject to constrains in Eq. (3)-(16). Here, $p_{i,m}$ indicates the quality of disinfection when completing task $i$ in mode $m$, $p^{max}$ is the highest possible quality of disinfection, and $\lambda$ represents the user-specified preference. When $\lambda=1$, the objective ignores quality and focuses on maximizing the number of tasks completed, whereas when $\lambda=0$, the objective exclusively prioritizes maximizing the quality of disinfection. This mixed integer optimization model can be solved using a standard branch and cut algorithm. However, for larger problem instances, the branch and cut solution has an exploding computation time. The CCG algorithm described in Alg. 1 can be used to effectively compute solutions in such cases.
(a) A sample layout for synthetic experiment.
(b) Comparison of idling time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(c) Comparison of working time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
(d) Impact on $SR$ and $DQ$ for different types of time windows. The tasks with time-windows could be critical or non-critical.
Figure 2: (a) A sample layout of a disinfection mission with different task types. (b) & (c) The solution of M3RS make effective use of agents as indicated by lower idling times and higher working times. (d) It can be seen that tasks available during the entire mission have a higher completion rate ($SR$) and disinfection quality as well.
IV Evaluation
In this section, we evaluate the the M3RS problem and propesed solution strategy of CCG. First, the metrics used for evaluation are defined. Second, the impact of using task-level quality as a decision variable is studied by comparing M3RS with the classic routing and scheduling in fixed (RS-F) disinfection mode on synthetically generated data. Then, the proposed CCG scheme is compared with a mixed integer programming (MIP) solver. Finally, hardware and simulated case studies are presented.
In each evaluation, multiple problem instances are generated. Each problem instance is characterized by mission time which is given by:
$T_{H}=\frac{t_{avg}\times\text{\# Tasks}}{\eta\times\text{Fleet size}}$ (24)
Here, $\eta$ is the load factor that is used to simulate missions with different $T_{H}$. A higher $\eta$ leads to tighter mission times. The average task completion time is given by $t_{avg}$. In our experiments, $t_{avg}$ is considered to be 4.5 minutes. We assume that the robot travels at a fixed speed of $0.5\,{m}/{s}$. We use the MIP solver from IBM’s CPLEX [33]. All the evaluations are carried out on machine equipped with an Intel i9 processor with 32 GB RAM.
IV-A Evaluation Metrics
We define some evaluation metrics to compare the performance of M3RS and RS-F (Sec. IV-B). The metrics are:
$SR\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}}{|\mathcal{T}|}$ (25)
$DQ\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}p_{i,m}\cdot y^{k}_{i,m}}{|\mathcal{T}|\cdot p^{max}}$ (26)
$MSI\,=\,\frac{SR+DQ}{2}$ (27)
$ATQ\,=\,\frac{DQ}{SR}$ (28)
$QBR\,=\,\frac{DQ-DQ^{*}}{(\bar{Q}_{exp}-\bar{Q}_{exp}^{*})+(\bar{M}_{exp}-\bar{M}_{exp}^{*})}$ (29)
The success rate ($SR$) measures the number of tasks completed successfully (Eq. (25)). The dosage quality ($DQ$) reflects the disinfection quality of completed tasks (Eq. (26)). The mean success index ($MSI$) is the average of $SR$ and $DQ$ (Eq. (27)). This metric scores a solution on combined performance on the number of tasks completed and disinfection quality. The average task quality ($ATQ$) measures the average disinfection quality received by a serviced task (Eq. (28)).
The quality benefit ratio ($QBR$) compares a target solution with a baseline solution in terms of the gains made in disinfection quality relative to the resource usage (Eq. (29)). It is a ratio of the difference in disinfection quality to the difference in resource expenditure, measured against a baseline solution. For the target solution, $\bar{Q}_{exp}$ and $\bar{M}_{exp}$ represent the fleet’s average energy expenditure and makespan. The RS-F solution serves as the baseline with $DQ^{*}$ as disinfection quality, $\bar{Q}_{exp}^{*}$ as average energy, and $\bar{M}_{exp}^{*}$ as makespan.
In Sec IV-C, we use relative optimal gap ($OG$) and the objective values (Eq. (23)) are also as a metric. The $OG$ measures the difference between the best known bound on the objective value ($o^{*}$) and the current solution’s objective value $o$, is also reported. MIP solvers typically provide this bound by solving a relaxed version of the problem and is given by:
$OG\,=\,\frac{o^{*}-o}{o^{*}}$ (30)
IV-B M3RS vs RS-F
IV-B1 Baselines
We assess the impact of the M3RS formulation, which incorporates quality levels as a decision variable through modes, compared to RS-F which is routing and scheduling with fixed execution modes. The RS-F formulation follows the vehicle routing problem with time windows [34], and we consider two variants: i) RS-F (Min), which uses the mode with the lowest disinfection quality, and ii) RS-F (Max), which uses the mode with the highest quality. For M3RS we use different values of user prefrence i.e. $\lambda$. A higher $\lambda$ prioritizes quantity over quality. Synthetically generated data is used to benchmark the problems, as detailed in the following paragraph. Both M3RS and RS-F (Min and Max) are solved using an MIP solver with a 10-minute time limit per problem instance. The evaluation metrics include $SR$, $DQ$, $MSI$ $ATQ$, and $QBR$, see Eq (25)-(29). RS-F (Min) is used as the baseline for $QBR$ calculation since it operates in the lowest disinfection mode, allowing us to measure the improvement in disinfection quality relative to resource expenditure. We analyze the solutions based on working and idle time and also discuss the impact of relaxing task time windows for M3RS.
IV-B2 Synthetic Data Generation
[…]
Table II: Experiment parameters for testing impact of quality.
Quantity Value
Layout area $150,200,225\,m^{2}$
Number of tasks $30-60$
Proportion of Critical tasks $15-25\%$
Avg. task completion time $4.5\,mins$
Task surface area range $0.5-2.2\,m^{2}$
$\eta$ 1
Fleet size 4
[…]
Table III: Comparison of performance metrics for synthetic problem instances. The colors indicate the first, second, and third ranking for a given metric. The solutions of M3RS perform well on both quality and quantity objectives as indicated by different metrics. The number of tasks range from $30-60$ and the robot fleet size is four.
Metric RS-F (Min) RS-F (Max) M3RS-0.01 M3RS-0.1 M3RS-0.5 M3RS-0.9
$SR$ $1.0\pm 0.0$ $0.847\pm 0.082$ $0.906\pm 0.078$ $0.912\pm 0.072$ $0.941\pm 0.068$ $0.973\pm 0.042$
$DQ$ $0.452\pm 0.031$ $0.757\pm 0.083$ $0.820\pm 0.078$ $0.829\pm 0.097$ $0.789\pm 0.095$ $0.750\pm 0.124$
$MSI$ $0.726\pm 0.016$ $0.802\pm 0.082$ $0.863\pm 0.072$ $0.870\pm 0.079$ $0.865\pm 0.073$ $0.862\pm 0.08$
$ATQ$ $0.452\pm 0.031$ $0.894\pm 0.024$ $0.906\pm 0.063$ $0.909\pm 0.070$ $0.839\pm 0.082$ $0.768\pm 0.104$
$QBR$ N/A $0.331\pm 0.064$ $0.364\pm 0.082$ $0.391\pm 0.117$ $0.331\pm 0.104$ $0.297\pm 0.129$
Table IV: Comparison of $OG$ values between MIP solver and CG solutions. Colors represent the first, second, and third rankings for the $OG$ metric. CCG solutions achieve better $OG$ values as compared to the MIP solver for the same computation time limit.
Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.9$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.9$
120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
MIP $0.131\pm 0.034$ $0.278\pm 0.113$ $0.247\pm 0.046$ $0.353\pm 0.024$ $0.253\pm 0.032$ $0.290\pm 0.061$ $0.570\pm 0.065$ $0.488\pm 0.039$
CCG ($K=2$) $0.136\pm 0.050$ $0.110\pm 0.024$ $0.218\pm 0.024$ $0.300\pm 0.018$ $0.116\pm 0.015$ $0.177\pm 0.039$ $0.455\pm 0.011$ $0.380\pm 0.006$
CCG ($K=4$) $0.126\pm 0.009$ $0.121\pm 0.034$ $0.211\pm 0.017$ $0.302\pm 0.023$ $0.097\pm 0.045$ $0.179\pm 0.021$ $0.434\pm 0.013$ $0.363\pm 0.014$
CCG ($K=6$) $0.110\pm 0.025$ $0.132\pm 0.025$ $0.208\pm 0.010$ $0.305\pm 0.021$ $0.107\pm 0.025$ $0.180\pm 0.021$ $0.453\pm 0.026$ $0.358\pm 0.029$
CCG ($K=8$) $0.133\pm 0.031$ $0.102\pm 0.022$ $0.220\pm 0.019$ $0.308\pm 0.009$ $0.106\pm 0.047$ $0.184\pm 0.020$ $0.446\pm 0.019$ $0.355\pm 0.025$
Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.1$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.1$
120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
MIP $0.124\pm 0.032$ $0.181\pm 0.021$ $0.153\pm 0.014$ $0.215\pm 0.019$ $0.409\pm 0.031$ $0.586\pm 0.013$ $0.5200\pm 0.026$ $0.602\pm 0.021$
CCG ($K=2$) $0.103\pm 0.018$ $0.144\pm 0.037$ $0.188\pm 0.022$ $0.205\pm 0.016$ $0.357\pm 0.037$ $0.483\pm 0.015$ $0.425\pm 0.018$ $0.514\pm 0.021$
CCG ($K=4$) $0.116\pm 0.018$ $0.126\pm 0.019$ $0.156\pm 0.029$ $0.197\pm 0.020$ $0.310\pm 0.023$ $0.491\pm 0.016$ $0.436\pm 0.021$ $0.517\pm 0.018$
CCG ($K=6$) $0.132\pm 0.027$ $0.150\pm 0.034$ $0.162\pm 0.032$ $0.201\pm 0.020$ $0.349\pm 0.031$ $0.483\pm 0.012$ $0.418\pm 0.009$ $0.520\pm 0.014$
CCG ($K=8$) $0.116\pm 0.019$ $0.134\pm 0.015$ $0.198\pm 0.010$ $0.181\pm 0.027$ $0.336\pm 0.011$ $0.491\pm 0.012$ $0.424\pm 0.015$ $0.502\pm 0.025$
IV-B3 Results
The mean and standard deviation of the performance metrics for solutions of RS-F and M3RS variants are presented in Table III. RS-F (Min) achieves the highest success rate ($SR$), addressing all tasks. This is because the lowest disinfection dosage is used by default, reducing disinfection time and increasing $SR$, but at the cost of lower disinfection quality ($DQ$). Conversely, RS-F (Max) provides higher $DQ$ but lower $SR$. The higher disinfection quality requires more time, limiting agents’ ability to complete other tasks. The solutions of M3RS outperform the RS-F baselines in the combined metric ($MSI$), offering a better balance between $SR$ and $DQ$. The solutions of M3RS admit better $DQ$ and $ATQ$ scores. The $QBR$ scores show that solutions of M3RS offer better quality improvement relative to the additional resource expenditure (energy and time). This suggests that considering disinfection quality through variable disinfection modes improves overall disinfection quality while maintaining or improving task completion rates.
We further analyze the missions by looking at the ratio of idle time to task servicing time, as well as the $SR$ and $DQ$ for different time-window types (see Fig. 2 (a)-(c)). The solution of M3RS show less idle time and more working time, indicating better use of mission time, see Fig. 2 (a)-(b) . Using variable disinfection modes likely saves time and energy, which can be spent on disinfecting more tasks or improving quality for others, depending on the user preference i.e. $\lambda$. RS-F (Max) shows higher idle time and lower working time than M3RS, as disinfection in the highest modes requires robots to arrive at or before a task’s time window. Early arrivals lead to waiting, limiting task completion and resulting in lower SR and DQ compared to M3RS.
When examining $SR$ and $DQ$ across time-window types, tasks available throughout the mission have the best performance, while those with limited time windows perform worse, see Fig. 2 (c). This suggests that relaxing time constraints where possible could lead to better overall performance.
IV-C Performance of CCG
IV-C1 Baselines
The performance of solving M3RS with CCG is compared with that of directly solving using MIP solver. For CCG, we report results with varying number of clusters i.e $K\in[2,4,6,8]$. The performance is evaluated in terms of the relative optimality gap ($OG$) given by Eq (30). For CCG, we use the best bound from the MIP solver to compute the $OG$.
IV-C2 Setup
For this, four problem types are considered with a number of tasks and robots mentioned in Table. IV. For each problem type, 5 instances are generated using the procedure outlined in Sec. IV-B. User preference for each instance is sampled from $\lambda\in[0.1,0.9]$, with higher $\lambda$ values prioritizing task completion over disinfection quality. We use load factors $\eta=1$ and $\eta=2$, where higher $\eta$ simulates missions with shorter $T_{H}$, providing less time for mission completion. The mission time ($T_{H}$) is calculated using Eq (24). The time limit for MIP solver and CCG is set to 10 minutes. We further report the performance with time limit for MIP set to 25 minutes.
IV-C3 Discussion
The mean and standard deviation of the $OG$ metric for MIP and CCG with time limit set to ten minutes is reported in Table IV. The change in $OG$ for MIP solver relative to CCG, with time limit for MIP solver set to 25 minutes is reported in Table V. The $OG$ achieved by CCG solutions is consistently better than that of MIP solver with the same time limit, see Table IV. Further, it is seen with increased time limit the performance of CCG remains comparable or better than MIP solver while being 2.5×2.5$\times$2.5 × faster. This indicates that clustering and sampling approach coupled with column generation provides an effective way to find good solution in shorter time as compared to MIP solver. Generally, clustering improves $OG$ values, with more clusters benefiting larger instances.
Table V: Difference of average optimality gap ($OG$) between CCG and the MIP solver across various instances. A 25-minute time limit is set for the MIP solver. Positive differences indicate better $OG$ for the MIP solver, while negative values favor CCG. Colors represent MIP outperforming CCG and CCG outperforming MIP. CCG achieves comparable or better $OG$ while being approximately 2.5×2.5$\times$2.5 × faster than the MIP baseline.
Metric $\eta$ $\lambda$ $120-4$ $150-5$ $175-5$ $200-6$
$OG_{CCG}-OG_{MIP}$ $1.0$ $0.9$ $0.038$ $-0.084$ $-0.058$ $-0.077$
$0.1$ $0.022$ $-0.052$ $-0.007$ $-0.026$
$2.0$ $0.9$ $0.003$ $-0.154$ $-0.074$ $-0.150$
$0.1$ $-0.029$ $-0.080$ $-0.091$ $-0.105$
IV-D Simulated and Hardware Case Study
(a) A team of G-robots disinfecting tables.
(b) The simulated environment with 20 tasks and 2 robots.
(c) The route sequences for simulated mission with 20 tasks, 2 robots, $T_{H}$ of 0.4 hours and $\lambda=0.9$.
(d) The route sequence for hardware experiment with $\lambda=0.9$
(e) The route sequence for hardware experiment with $\lambda=0.1$
Figure 3: (a) A team of G-robots [32] disinfecting tables. (b) The routes of two robots executing various tasks in the simulated environment. The route of the first robot is highlighted in white and for the second robot yellow is used. (c) The route sequence taken by each robot in the simulated mission. For $\lambda=0.9$, maximization of the number of tasks completed has more emphasis than the quality of disinfection. As a result, most of the tasks are done in higher modes. (d) & (e) The route sequence taken by each robot for the hardware mission with $\lambda=0.9$ and $0.1$. For $\lambda=0.9$, the emphasis is on disinfecting all the tables at higher disinfection modes. Whereas, for $\lambda=0.1$, the emphasis is on ensuring high-quality disinfection rather than disinfecting all tables.
IV-D1 Simulation and Hardware Experiment Setup
Multiple disinfection mission instances are generated for a simulated environment. Each disinfection mission is characterized by mission time, time windows for each task, and number of robots used in the mission. The simulated environment consists of $20$ tables as disinfection surfaces spread across two rooms. The fleet size considered in the range, $|\mathcal{A}|=2-5$. The mission time is calculated using Eq. 24. The entire simulated experiment is conducted in ROS [35] and Gazebo environment [36]. A simulated mission is depicted in Fig 3 (b).
For the hardware case study, we use two G-robots (shown in Fig. 3 (a)) for a disinfection mission consisting of six tables. Each table can be disinfected in three disinfection modes, i.e. $D_{99.99}$, $D_{99}$, and $D_{90}$. The entire disinfection mission must be completed in $17$ minutes.
For both experiments, M3RS is solved for two preferences: $\lambda\,=0.9$ and $0.1$ for each mission. The optimization model is solved in CPLEX [33]. The metrics reported include $SR$, $DQ$, $MSI$, and average lateness of the robot in task completion.
IV-D2 Discussion
The metrics for simulated and hardware case studies are reported in Table VI. The example of generated route sequences in simulated experiment is visualized in Fig. 3 (b)-(c). The route sequence for hardware experiments is shown in Fig. 3 (d)-(e). From Table VI, it can be seen that a decent $SR$ is achieved in all missions. Even for missions with tight $T_{H}$ the $SR$ is high. Therefore, a larger robotic fleet would be necessary to compensate for tighter mission times in order to service as many tasks as possible. For missions with shorter $T_{H}$, there is a decline in $DQ$ metric. This is because with tighter mission times, each task has a smaller time window, which means many of the tasks cannot be performed in higher disinfection modes. A possible remedy could be to enable robots to collaborate on different tasks. Finally, there are some minor delays observed in terms of task completion times for simulation and hardware implementation. These delays are caused by slightly longer path navigation times. Such delays can be accounted for by introducing chance constraints to model stochasticity in the system.
Table VI: Results for simulated and hardware experiments.
Num. of tasks and robots $T_{H}$ (hrs) $\lambda$ $SR$ $DQ$ MSI Avg. Lateness (s)
20-2 0.40 0.9 0.80 0.57 0.63 1.9
0.1 0.70 0.71 0.60 6.2
20-3 0.21 0.9 0.80 0.51 0.60 30.1
0.1 0.75 0.56 0.58 1.4
20-4 0.20 0.9 0.95 0.55 0.74 33.6
0.1 0.90 0.60 0.72 19.1
20-5 0.16 0.9 0.80 0.57 0.68 29.4
0.1 0.75 0.62 0.68 33.7
6-2* 0.28 0.9 1.00 0.40 0.70 27.0
0.1 0.67 1.00 0.65 32.0
* = Results for hardware experiments.
V Conclusions and Future work
In this work, we proposed M3RS, a novel multi-objective formulation that routes and schedules a team of agents to accomplish various tasks in a mission. The agents can perform the task in one of many available modes. Each mode has distinct task quality, resource consumption, and service times. We first propose a mixed-integer linear programming model for M3RS. Then, a column generation approach is presented to tackle larger instances with a faster computation time. We showcase the utility of M3RS in multi-robot disinfection applications. The experiments suggest that our solution for M3RS provides more flexibility to users to select solutions based on desired preferences. The current approach is limited as it assumes a deterministic environment, unlike real-world scenarios. Generating the entire Pareto front is computationally expensive, particularly for larger instances. Our future work will extend M3RS to incorporate elements like stochastic travel times, dynamic arrival of tasks and realistic issues like robot breakdown. Further, we aim to tightly integrate the robot path planning with the decision making framework provided by M3RS. Finally, we will investigate using deep reinforcement learning techniques to generate good quality solutions in real-time.
Acknowledgements
We would like to thank Mr. Matthew Lisondra for all the support with the experiments.

### v3

@@
 Abstract
-The quality of task execution can significantly impact a multi-robot mission.
-While higher quality is desirable, it may not always be feasible due to mission constraints.
-Existing multi-robot task allocation literature generally overlooks quality of service as a decision variable.
-Addressing this gap, we introduce the multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem, designed for time-bound, multi-robot, multi-objective missions.
-In M3RS, each task offers multiple execution modes, each with different resource requirements, execution time, and quality.
-M3RS optimizes task sequences and execution modes for each agent.
-The need for M3RS comes from multi-robot applications in which a trade-off between multiple criteria can be achieved by varying the task level quality of service through task execution modes.
-Such ability is particularly useful for service robot applications.
-We use M3RS for the application of multi-robot disinfection in healthcare environments and other public locations.
-The objectives considered for disinfection application are disinfection quality and number of tasks completed.
-A mixed-integer linear programming (MIP) model is proposed for M3RS.
-Further, a clustering-based column generation (CCG) algorithm is proposed to handle larger problem instances.
-Through synthetic, simulated, and hardware case studies, we demonstrate the advantages of M3RS, showing it provides flexibility and strong performance across multiple metrics.
-Our CCG algorithm generates solutions 2.5$\times$ faster than a baseline MIP optimizer, maintaining competitive performance.
-The videos for the experiments are available on the project website [^1] https://sites.google.com/view/g-robot/m3rs/.
-Index Terms:
-multi-robot system, autonomous systems, task allocation, scheduling, service robots, disinfection robots.
+[…]
 I Introduction and Related Works
-Maintaining hygiene in healthcare environments is crucial to prevent catastrophic issues like hospital-acquired infections.
-Disinfection robots are useful technologies that provide an additional layer of protection by automating the disinfection of target surfaces.
-A team of these robots can improve both spatial and temporal efficiency in maintaining hygiene standards in hospitals [1].
-To coordinate a fleet of robots for a disinfection mission, two key decisions must be made: i) which task will be assigned to which robot, and ii) what level of disinfection quality should be applied to each task.
-While higher disinfection quality is desirable, it also requires more time and resources, which may not always be feasible.
-The assignment of tasks to robots has been studied in the literature [2], but there is a lack of research addressing the decision regarding the quality of service of a task.
-In disinfection applications, considering the quality of tasks is important as it impacts overall hygiene standards.
-Additionally, some surfaces, such as those in patient rooms and ICUs, require stringent quality standards.
-In this work, we tackle the problem of deciding which tasks to implement and at what mode/quality.
-For a team of robots, Multi-robot Task Allocation (MRTA) algorithms [3, 2] allocate tasks to various robots working on a common mission while optimizing a joint mission objective.
-A key metric to evaluate these allocations is the quality of service (QoS) for the mission.
-The definition of QoS depends on the mission requirements and target application, such as the quality of communication between the robot and the base station [4, 5], the tardiness of addressing tasks [6], the average waiting time for a customer [7], or productivity and fatigue [8].
-Often, there is a trade-off between QoS and other mission objectives, which could be conflicting.
-In such cases, multiple solutions exist that prioritize these objectives [4, 5, 8].
-Figure 1: M3RS is suitable for routing and scheduling in time-bound scenarios where multiple criteria of interest depend on the mode in which the agents execute the tasks.
-This visualization depicts two cases.
-Here, mode 1 stands for the highest quality and long task execution times, whereas mode 3 is for the lowest quality and fastest task execution times.
-The left side of the image shows a case where agents choose to maximize the number of tasks completed by performing tasks in lower-quality modes.
-On the right, agents prioritize quality over quantity.
-As a result, tasks 2 and 5 are missed.
-An example of such a scenario is a multi-robot system performing services such as disinfection and cleaning.
-MRTA problems are $\mathcal{NP}$-hard, making larger instances challenging to solve.
-With extra variables, concerning the quality of the service, the problem becomes more complex.
-Techniques such as heuristics and metaheuristics are faster but often sacrifice optimality [9].
-Learning-based methods require expensive training and massive datasets, and may still struggle to generalize [10].
-Column generation is a decomposition approach that refines the solution iteratively by solving smaller sub-problems.
-This allows faster convergence to higher-quality solutions as it directly exploits the problem’s structure and efficiently handles large-scale instances.
-We propose a clustering-based column generation algorithm, where clustering further helps in reducing the search space to find good quality solutions.
-To the best of our knowledge, this is the first work to address task-level service quality using multiple execution modes for quality-dependent multi-robot missions.
-We introduce a multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem.
-The contributions of this work are:
+Multi-robot teams are widely used in diverse applications such as search and rescue (Liu and Nejat,, 2016; Olson et al.,, 2012), environment mapping (Saeedi et al.,, 2016; Lajoie et al.,, 2022), disaster management response (Phan and Liu,, 2008), marine operations (Thompson and Guihen,, 2019), and agricultural tasks (Das et al.,, 2025).
+In healthcare settings, these teams can play a vital role in maintaining hygiene to prevent hospital-acquired infections.
+Disinfection robots offer an effective solution by automating surface disinfection, adding an extra layer of protection.
+A team of these robots can enhance both spatial and temporal efficiency in maintaining hygiene standards in hospitals (Mehta et al.,, 2023).
+Coordinating a fleet of disinfection robots requires making two key decisions: i) assigning tasks to robots, and ii) determining the level of disinfection quality for each task.
+Although higher disinfection quality is preferable, it also requires more time and resources, which may not always be feasible.
+Task allocation for multi-robot systems has been well-studied in the literature (Korsah et al.,, 2013), but the decision regarding quality of service (QoS) for individual tasks remains largely unexplored.
+In disinfection applications, QoS is critical, as some surfaces, such as those in patient rooms and ICUs, require higher hygiene standards.
+This work addresses the challenge of selecting tasks and determining the appropriate disinfection quality/mode for each task.
+Multi-Robot Task Allocation (MRTA) algorithms (Gerkey and Matarić,, 2004; Korsah et al.,, 2013) assign tasks to robots to optimize a joint mission objective.
+A key factor in evaluating these allocations is the QoS of the mission.
+The target of QoS varies depending on the application.
+For instance, QoS can refer to the quality of communication between robots and a base station (Khan et al.,, 2022; Wang et al.,, 2023), reduced task completion delays (Botros et al.,, 2023), quantity of area coverage for agricultural survey applications (Barrientos et al.,, 2011), reduced average customer waiting time (Miller and How,, 2017), or optimal worker productivity and reduced fatigue (Calzavara et al.,, 2023).
+In many cases, QoS competes with other mission objectives, requiring trade-offs among different priorities (Hayat et al.,, 2020; Flushing et al.,, 2017; Calzavara et al.,, 2023; Liang et al.,, 2023).
+While QoS has been considered to some extent in the MRTA literature, no prior work incorporates QoS as a decision variable specifically for disinfection applications.
+Introducing QoS as an explicit decision variable enables more precise control over the trade-offs between disinfection quality and other mission objectives.
+In disinfection scenarios, for example, it allows for optimizing hygiene standards while maximizing task completion within the available mission time.
+Agent 1Agent 2Task 1Task 2Task 4Task 3(a)Agent 1Agent 2Task 1Task 2Task 4Task 3(b)Agent 1Agent 2Task 1Task 2Task 4Task 3(c)Mode 1Mode 2Mode 3Missed Task
+Figure 1: Traditional Multi-Robot Task Allocation models overlook Quality of Service (QoS) as a decision variable, assuming agents provide a uniform service level, which can lead to inefficiencies in routing and scheduling under time and resource constraints.
+The visualization compares three cases: (a) Agents operate at the highest service level (Mode 1), ensuring maximum quality but consuming more time and energy, resulting in missed tasks (e.g., Tasks 2 and 3).
+(b) Agents operate at the lowest service level (Mode 3), completing all tasks but at reduced quality.
+(c) M3RS incorporates QoS as a decision variable, allowing agents to balance quality and quantity dynamically–choosing from Mode 1, Mode 2 (moderate quality), and Mode 3–ensuring all tasks are addressed at appropriate service levels.
+Unlike conventional methods, M3RS enables flexible task allocation by integrating QoS into decision-making, making it well-suited for multi-robot applications such as disinfection and cleaning, where balancing quality and quantity is crucial.
+Additional decision variables for QoS can further increase the computation complexity of the MRTA problem, which is inherently $\mathcal{NP}$-hard.
+While heuristics and metaheuristics can provide faster solutions, they often compromise optimality (Braekers et al., 2016b, ).
+Learning-based approaches require extensive training data and may struggle to generalize (Bogyrbayeva et al.,, 2022).
+To efficiently handle the added complexity of task-level QoS variables, we propose a clustering-based column generation algorithm.
+Column generation (Desrosiers and Lübbecke,, 2005) iteratively refines solutions by solving smaller subproblems, allowing for faster convergence to high-quality solutions.
+The integration of clustering further helps in reducing the search space to find good quality solutions.
+To the best of our knowledge, this work presents the first MRTA problem to incorporate task-level QoS using multiple execution modes for quality-dependent multi-robot missions.
+We introduce the Multi-Robot, Multi-Objective, and Multi-Mode Routing and Scheduling (M3RS) problem, which enables adjustment of task-level QoS to optimize trade-offs among mission objectives.
+M3RS is particularly relevant for applications where tasks are spatially distributed, have varying service and resource requirements, and involve multiple mission criteria.
+Each robot can execute tasks in different modes, where each mode dictates task quality, resource consumption, and service time.
+A visual representation of M3RS compared to the typical MRTA problem is provided in Fig.
+1.
+Traditional MRTA models ignore QoS, while we consider it as a decision variable.
+We use M3RS for multi-robot disinfection, where both disinfection quality and task throughput are critical.
+High-quality disinfection, however, increases time and energy costs, limiting the number of tasks performed.
+M3RS addresses this trade-off by leveraging task-execution modes to balance quality and efficiency.
+Our contributions are briefly summarized below:
   1. 1.
-    We present M3RS, a novel problem to alter task-level QoS to achieve trade-offs in different mission objectives.
-The proposed problem is well-suited for applications where tasks are spread across a physical space, each task has service requirements and resource requirements, and the overall mission has multiple criteria of interest.
-Each robot can address the task in multiple execution modes, where each mode corresponds to different task quality, resource consumption, and service times.
-A visual depiction of the application of M3RS is given in Fig.
-1.
+    We propose a novel MRTA problem called Multi-Robot, Multi-Objective, and Multi-Mode Routing and Scheduling (M3RS) that explicitly considers task-level QoS.
   2. 2.
-    We present the usage of M3RS for a multi-robot disinfection application.
-A thorough study of different mission scenarios is conducted on a synthetic dataset.
-Further, we test M3RS in simulated and hardware case studies.
+[…]
   3. 3.
-    We use a mixed-integer linear programming formulation for the proposed problem.
-To handle the large decision space, we propose clustering-based column generation (CCG) for solving M3RS.
-CCG generates solutions 2.5$\times$ faster than the baseline while maintaining competitive performance.
+[…]
+  4. 4.
+    We present simulated and real-world experiments for multi-robot disinfection missions.
 This paper is organized as follows: The general formulation and optimization of M3RS is given in Section III.
@@
 Finally, Section V includes future works and conclusions.
-II Literature Review: QoS in Task Assignemnt
-QoS has also received attention in routing and production scheduling literature.
-Exposito et al.
-[11] considered the timeliness of vehicles as a QoS indicator for vehicle routing.
-Similarly, QoS has been studied for home healthcare applications, where it involves routing nurses to different patients for healthcare-related services, depending on the patient receiving care from a nurse with the desired experience [12, 13].
-In supply chain and production planning, QoS has been considered as the existence of minimum stock [14].
+II Literature Review: QoS in Task Assignment
+QoS has also received attention in the literature on routing and production scheduling.
+Expósito et al., (2019) considered the timeliness of vehicles as a QoS indicator for vehicle routing.
+Similarly, QoS has been studied for home healthcare applications, where it involves routing nurses to different patients for healthcare-related services, depending on the patient receiving care from a nurse with the desired experience (Kordi et al.,, 2023; Di Mascolo et al.,, 2021).
+In supply chain and production planning, QoS has been considered as the existence of minimum stock (Anli et al.,, 2007).
+In logistics application, customer satisfaction (Liang et al.,, 2023) and optimizing the service times (Gonzalez-R et al.,, 2024) have also been considered as indicators of QoS.
 Most of these existing works modify the task allocation to achieve different trade-offs between QoS and other objectives.
-However, there are no works that focus on the deliberate degradation or improvement of QoS offered to an individual task by a robot as a means to achieve an overall trade-off for various mission objectives.
-Task execution modes are different operational settings for servicing tasks.
-These modes depend on various constraints in the environment or aim to achieve desired performance on key metrics like quality and resource consumption.
-Lahijanian et al.
-[15] explored using an energy-saving mode where resource-intensive autonomy features are disabled to conserve energy in resource-constrained environments.
-The usage of task execution modes in MRTA has been limited.
-Task execution modes were used to decide if humans and robots should collaborate to maximize time efficiency [16, 17].
-In project scheduling, task execution modes are used to minimize project makespan while adhering to resource constraints [18, 19, 20].
-In vehicle routing, mode selection involves choosing delivery locations with specific constraints, impacting customer delivery [21].
-The hybrid electric vehicle routing problem in [22] used mode selection for switching driving modes, affecting emissions and transportation costs.
-Using task execution modes to select QoS provided to the task is still unexplored.
-This decision affects the quality of individual tasks and the overall mission.
-For example, in a disinfection mission, cleaning a surface with the highest available dosage ensures the highest quality level.
-However, higher dosages require longer service times, which could potentially result in skipping other tasks.
+However, there are no works that focus on the deliberate degradation or improvement of QoS offered to an individual task as a means to achieve an overall trade-off for various mission objectives.
+Task execution modes refer to distinct operational settings for completing tasks, often chosen based on environmental constraints or desired performance metrics such as tardiness and resource consumption.
+This concept has been extensively studied in domains like project scheduling and vehicle routing.
+In project scheduling, execution modes typically represent alternative ways to perform an activity—for instance, allocating more resources to reduce its duration (Hartmann and Briskorn,, 2022; Wkeglarz et al.,, 2011; Tiwari et al.,, 2009).
+In routing applications, mode selection has been used to choose between transportation methods for healthcare service providers (Braekers et al., 2016a, ), manage hybrid vehicle driving strategies to minimize emissions and costs (Seyfi et al.,, 2022), and optimize battery recharging strategies in electric vehicle fleets (Zhang et al.,, 2024).
+In MRTA, the use of task execution modes has been limited.
+Some studies have explored their use to determine whether human-robot collaboration is beneficial for improving time efficiency (Liau and Ryu,, 2022; Ferreira et al.,, 2021; Parreño-Torres et al.,, 2024).
+However, the use of task execution modes to explicitly control the QoS provided to individual tasks remains largely unexplored in robotic applications.
+This is particularly important in domains such as multi-robot disinfection, where higher disinfection quality (achieved through higher dosages, for example), has clear physical significance but also incurs greater time and energy costs, potentially reducing overall task throughput.
+We use task execution modes in M3RS to explore such trade-offs in the context of MRTA.
 III Multi-Robot Multi-Objective Multi-Mode Routing and Scheduling
 In this section, first, the general optimization model for M3RS problem is presented.
-Then we present a column generation (CG) scheme to solve the optimization model with faster computation times.
+Then, we present a column generation (CG) scheme to solve the optimization model with faster computation times.
 Our novel contribution is the use of clustering heuristics, which helps solve the problem more quickly and find solutions with better objective values.
@@
 Table I: Notation for M3RS.
-Notation Description
-$\mathcal{A}$ Set of agents.
-$\mathcal{T}$ Set of tasks.
-$\mathcal{M}_{i}$ Set of modes for task $i$, where $\mathcal{M}_{i}=\{1,2...M\}$.
-$R$ Number of objective criteria.
-$f_{r}$ $r^{th}$ objective criterion.
-$p_{i,m}$ The quality indicator of doing task $i$ in mode $m$.
-$p^{max}$ The highest possible quality of disinfection
-$\sigma_{i,m}$ Time to complete task $i$ in mode $m$.
-$q_{i\scriptsize{\xrightarrow{}}j}$ Energy consumed for travelling from task $i$ to task $j$.
-$q_{idle}^{k}$ Energy consumption during idling by agent $k$.
-$q_{i,m}$ Energy consumption of task $i$ in mode $m$.
-$\text{q}_{\text{travel}}$ Avg. current consumption in traveling at a constant speed.
-$\text{q}_{\text{tc}}$ Avg. current consumption while completing tasks.
-$\text{q}_{\text{idling}}$ Avg. current consumption while idling.
-$[a_{i},b_{i}]$ Start and end time for task $i$.
-$Q$ Total capacity of an agent.
-$T_{H}$ Time available to complete all tasks.
-$t_{i,j}$ Travel time from task $i$ to task $j$.
-$\lambda_{r}$ The scalar preference for $r^{th}$ objective function.
-$s^{k}_{i}$ A decision variable for arrival time of agent $k$ to task $i$.
-$x^{k}_{i{\xrightarrow{}}j}$ A binary decision variable for if agent $k$ goes from task $i$ to task $j$.
-$y^{k}_{i,m}$ A binary decision variable for task $i$ done in mode $m$ by agent $k$.
-$\eta$ The load factor hat is used to simulate different disinfection missions.
+No Notation Description Comment
+1 $\mathcal{A}$ Set of agents. Problem Sets
+2 $\mathcal{T}$ Set of tasks.
+3 $\mathcal{M}_{i}$ Set of modes for task $i$, where $\mathcal{M}_{i}=\{1,2...M\}$.
+4 $R$ Number of objective criteria. Objective Function
+5 $f_{r}$ $r^{th}$ objective criterion.
+6 $\lambda_{r}$ The scalar preference for $r^{th}$ objective Function.
+7 $q_{i\scriptsize{\xrightarrow{}}j}$ Energy consumed for traveling from task $i$ to task $j$. Energy Constraints
+8 $q_{\text{idle}}^{k}$ Energy consumption during idling by agent $k$.
+9 $q_{i,m}$ Energy consumption of task $i$ in mode $m$.
+10 $q_{\text{travel}}$ Average current consumption in traveling at a constant speed.
+11 $q_{\text{tc}}$ Average current consumption while completing tasks.
+12 $q_{\text{idling}}$ Average current consumption while idling.
+13 $Q$ Total energy capacity of an agent.
+14 $\sigma_{i,m}$ Time to complete task $i$ in mode $m$. Time Constraints
+15 $a_{i},b_{i}]$ Start and end time for task $i$.
+16 $T_{H}$ Mission time or time available to complete all tasks.
+17 $t_{i,j}$ Travel time from task $i$ to task $j$.
+18 $p_{i,m}$ The quality reward of doing task $i$ in mode $m$. Disinfection Problem
+19 $p^{max}$ The reward for the highest possible quality of disinfection.
+20 $\eta$ Load factor: mission workload relative to the robot’s available capacity.
+21 $x^{k}_{i{\xrightarrow{}}j}$ A binary decision variable for if agent $k$ goes from task $i$ to task $j$. Decision Variables
+22 $y^{k}_{i,m}$ A binary decision variable for task $i$ done in mode $m$ by agent $k$.
+23 $s^{k}_{i}$ A decision variable for arrival time of agent $k$ to task $i$.
 III-A Optimization Model
-A solution for the M3RS problem determines routes, schedules, and task execution modes for a set of agents $\mathcal{A}$, and a set of tasks $\mathcal{T}$ distributed across physical space.
-Refer to Table I for notation.
-All the agents are identical, and only one agent can service each task.
+A solution for the M3RS problem determines routes, schedules, and task execution modes for a set of agents $\mathcal{A}$, and a set of tasks $\mathcal{T}$, distributed across a physical space.
+Refer to Table I for notation and symbols used to formulate the problem.
+Here, all agents are identical, and only one agent can service each task.
 We use $i$ and $j$ to denote different tasks in $\mathcal{T}$ and $k$ to denote $k^{th}$ agent in the rest of the notation.
@@
 Performing a task in a mode with higher service time and resource consumption would lead to lower availability of agents for other tasks.
-The overall scheduling and allocation mission is optimized for multiple objectives.
-The energy capacity for each agent is given by $Q$ Ah (Ampere-hours).
-The energy consumption for each agent is divided into performing tasks, traveling, and idling [23].
-A fixed-rate battery model is used to calculate the energy consumption, where energy consumed is proportional to idling time, task execution time, and traveling time respectively [23].
-Different aspects of energy expenditure can be calculated using:
-$\begin{split}q_{i\scriptsize{\xrightarrow{}}j}=\text{q}_{\text{travel}}\cdot t_{i,j},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e.
-traveling energy)}\\ q_{i,m}=\text{q}_{\text{tc}}\cdot\sigma_{i,m},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e.
-servicing energy)}\\ q_{idle}^{k}=\text{q}_{\text{idle}}\cdot(M_{k}-W_{k}).\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e.
-idling energy)}\end{split}$ (1)
-Here, $\text{q}_{\text{travel}}$ is the current consumption for an agent traveling at a constant speed, $\text{q}_{\text{tc}}$ is the current consumption while performing a task, and $\text{q}_{\text{idle}}$ is the current consumption for idling.
-$M_{k}$ is the total time spent by agent $k$ on the mission, and $W_{k}$ is the total time spent by agent $k$ traveling and executing tasks.
-The net energy consumed by agent $k$ in a mission is given by the LHS of Eq.
-(9).
-The relevant notation for the formulation is summarised in Table I.
+The overall scheduling and allocation mission is optimized for $R$ objective functions.
 The formulation for M3RS is expressed as a mixed integer with the following decision variables:
-i) the binary decision variable $x^{k}_{i{\xrightarrow{}}j}\in\{0,1\}$ for assignment and transition of agent $k$ from task $i$ to $j$;
-ii) the binary decision variable $y^{k}_{i,m}\in\{0,1\}$ for allocation of mode $m$ for task $i$ when executed by agent $k$; and
-iii) the real continuous variable $s^{k}_{i}\in\mathbb{R}$ denotes arrival time of agent $k$ to task $i$.
-The optimization problem is expressed as
+  1. 1.
+    The binary decision variable $x^{k}_{i{\xrightarrow{}}j}\in\{0,1\}$, for assignment and transition of agent $k$ from task $i$ to $j$;
+  2. 2.
+    The binary decision variable $y^{k}_{i,m}\in\{0,1\}$, for allocation of mode $m$ for task $i$ when executed by agent $k$; and
+  3. 3.
+    The real continuous variable $s^{k}_{i}\in\mathbb{R}$ denotes arrival time of agent $k$ to task $i$.
+With these descriptions, the optimization problem is expressed as follows, optimizing the variables mentioned above:
 $\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m},s^{k}_{i}}{\max}\,\sum_{r=0}^{R}\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m},s^{k}_{i}),$ (1)
@@
 (1) is the scalar objective function that maximizes the convex combination of $R$ objective criteria.
-Here, $f_{r}$ is $r^{th}$ objective criteria for $r=1,2,...R$.
-The user-specified parameter $\lambda_{r}$ represents trade-off preference for $r^{th}$ objective criterion.
+Here, $f_{r}$ is $r^{th}$ objective criteria for r=1,2,Rr=1,2,...R.
+The user-specified parameter $\lambda_{r}$ represents the trade-off preference for $r^{th}$ objective criterion.
 Each objective criterion can depend on task allocation, selected mode, and arrival times.
-The constraints of the optimization are categorized into five sets as follows: 1) task assignment constraint, 2) feasibility constraints, 3) energy consumption constraint, 4) time constraints, and 5) domain and preference constraints.
+The constraints of the optimization, described through Eq. (2)-(16), are categorized into five categories, as follows:
+  1. 1.
+    Task assignment constraint, Eq. (2),
+  2. 2.
+    Feasibility constraints, Eq. (3)-(7),
+  3. 3.
+    Energy consumption constraint, Eq. (9),
+  4. 4.
+    Time constraints, and Eq. (10)-(12), and
+  5. 5.
+    Domain and preference constraints, Eq. (13)-(16).
 These constraints are described next.
@@
 Given by Eq. (2), the task assignment constraint ensures that each task is done at most once:
-[…]
+$\sum_{k\in\mathcal{A}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}\,\leq\,1\,\,\,\forall i\in\mathcal{T}\backslash 0.$ (2)
 Here, $i=0$ is excluded ($\mathcal{T}\backslash 0$) as it corresponds to the home depot.
-This is done to maintain feasibility as all agents must leave the home depot.
+This is done to maintain feasibility, as all agents must leave the home depot.
 There might be missions where some tasks must be completed.
-For such cases, the inequality can be replaced by equality for tasks that must be completed.
+For such cases, where the tasks must be completed, the inequality can be replaced by equality.
 III-A2 Feasibility Constraints
 Given by Eq. (3)-(7), the feasibility constraints ensure the solutions are logical.
-[…]
-[…]
-[…]
-[…]
-[…]
+$x^{k}_{i{\xrightarrow{}}i}\,=\,0\,\,\forall k\in\mathcal{A}\,\,\forall i\in\mathcal{T},$ (3)
+$\sum_{i\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}0}\,=\,1\,\,\,\forall k\in\mathcal{A},$ (4)
+$\sum_{j\in\mathcal{T}}x^{k}_{0\scriptsize{\xrightarrow{}}j}\,=\,1\,\,\,\forall k\in\mathcal{A},$ (5)
+$\sum_{i\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}h}-\sum_{j\in\mathcal{T}}x^{k}_{h\scriptsize{\xrightarrow{}}j}\,=\,0\,\,\,\forall h\in\mathcal{T},\,\,\forall k\in\mathcal{A},$ (6)
+$\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}-\sum_{m\in\mathcal{M}_{i}}y^{k}_{i,m}\,=\,0\,\,\forall(k,i)\in\mathcal{A}\cup\mathcal{T}\backslash 0.$ (7)
 The constraint in Eq.
@@
 III-A3 Energy Consumption Constraint
+The energy capacity for each agent is given by $Q$ Ah (Ampere-hours).
+The energy consumption for each agent is divided into performing tasks, traveling, and idling (Zhang et al.,, 2023).
+A fixed-rate battery model is used to calculate the energy consumption, where energy consumed is proportional to idling time, task execution time, and traveling time, respectively (Zhang et al.,, 2023).
+Different aspects of energy expenditure can be calculated using:
+$\begin{split}q_{i\scriptsize{\xrightarrow{}}j}=q_{\text{travel}}\cdot t_{i,j},\penalty 10000\ \penalty 10000\ \text{\small(i.e.
+traveling energy)}\\ q_{i,m}=q_{\text{tc}}\cdot\sigma_{i,m},\penalty 10000\ \penalty 10000\ \text{\small(i.e.
+servicing energy)}\\ q_{\text{idle}}^{k}=q_{\text{idling}}\cdot(M_{k}-W_{k}).\penalty 10000\ \penalty 10000\ \text{\small(i.e.
+idling energy)}\end{split}$ (8)
+Here, $q_{\text{travel}}$ is the current consumption for an agent traveling at a constant speed, $q_{\text{tc}}$ is the current consumption while performing a task, and $q_{\text{idling}}$ is the current consumption for idling.
+$M_{k}$ is the total time spent by agent $k$ on the mission, and $W_{k}$ is the total time spent by agent $k$ traveling and executing tasks.
+The net energy consumed by agent $k$ in a mission is given by the LHS of Eq.
+(9).
+The relevant notation for the formulation is summarized in Table I.
 As described by Eq. (9), the energy consumption constraint enforces that the energy expenditure of an agent in the form of traveling, doing tasks, and idling does not exceed the agent’s capacity:
-$\text{q}_{\text{idle}}^{k}+\sum_{i\in\mathcal{T}}(\sum_{m\in\mathcal{M}_{i}}q_{i,m}\cdot y^{k}_{i,m}+\sum_{j\in\mathcal{T}}q_{i\scriptsize{\xrightarrow{}}j}\cdot x^{k}_{i{\xrightarrow{}}j})\,\leq\,Q\,\,\,\forall k\in\mathcal{A}.$ (9)
+$q_{\text{idle}}^{k}+\sum_{i\in\mathcal{T}}(\sum_{m\in\mathcal{M}_{i}}q_{i,m}\cdot y^{k}_{i,m}+\sum_{j\in\mathcal{T}}q_{i\scriptsize{\xrightarrow{}}j}\cdot x^{k}_{i{\xrightarrow{}}j})\,\leq\,Q\,\,\,\forall k\in\mathcal{A}.$ (9)
 III-A4 Time Constraints
-Given by Eq. (10)-(12), the time constraints ensure the timing of the tasks are consistent:
-$s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}+t_{i,j}-s^{k}_{j}\leq 0\,\,\forall i\in\mathcal{T}$ (10)
-$\forall j\in\mathcal{T}\,\,\forall k\in\mathcal{A},$
+Given by Eq. (10)-(12), the time constraints ensure the timing of the tasks is consistent:
+$s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}+t_{i,j}-s^{k}_{j}\leq 0\,\,\forall i\in\mathcal{T}\\ \forall j\in\mathcal{T}\,\,\forall k\in\mathcal{A},$ (10)
 $a_{i}\leq s^{k}_{i}+\sum_{m\in\mathcal{M}_{i}}\sigma_{i,m}\cdot y^{k}_{i,m}<b_{i}\,\,\,\forall i\in\mathcal{T},\,\,\forall k\in\mathcal{A},$ (11)
@@
 (10) ensures that the arrival time for a subsequent task is always later than the previous task.
-This constraint also enables sub-tour elimination [24], where sub-tours are infeasible solutions that represent disjoint sequences of completing tasks for the same agent.
+This constraint also enables sub-tour elimination (Kara et al.,, 2004), where sub-tours are infeasible solutions that represent disjoint sequences of completing tasks for the same agent.
 The constraint in Eq.
@@
 III-B A Solution for M3RS: Clustering-based Column Generation
-M3RS extends traditional routing and scheduling problems by adding the decision of mode selection, which impacts task quality.
-Routing and scheduling problems are known to be $\mathcal{NP}-$hard [25].
-Consequently, larger problem instances with many tasks and agents lead to exponential growth in the solution space and significantly burden computational resources [26].
-We use CG [27, 28], to address the computational challenges associated with the large decision variables in the proposed formulation.
-The strength of CG lies in its ability to break down a complex problem into manageable sub-problems while iteratively refining the solution.
-The CG algorithm decomposes the original optimization problem into a master problem and a sub-problem.
-The master problem identifies optimal columns (or paths/solutions) from a pool of feasible candidate solutions, $\Omega$.
-Since $\Omega$ is very large, typically, a restricted master problem is solved on a subset $\Omega^{\prime}\subset\Omega$.
-The sub-problem uses the dual solution from the master problem to find new favorable columns, which are added to $\Omega^{\prime}$ to improve the objective value of the master problem.
-The master problem and sub-problem are solved iteratively until there are no new columns with positive reduced cost.
-We refer to our approach as CCG (clustering-based CG) due to the use of clustering-based heuristics for solving the sub-problem.
-The rest of the section presents the building blocks of our CCG algorithm.
+M3RS extends traditional routing and scheduling problems by incorporating the decision of mode selection, which affects task quality.
+Routing and scheduling problems are known to be $\mathcal{NP}$-hard (Kumar and Panneerselvam,, 2012).
+Consequently, larger problem instances with many tasks and agents lead to exponential growth in the solution space and significantly burden computational resources (Uchoa et al.,, 2017).
+We use Column Generation (CG) (Desrosiers and Lübbecke,, 2005; Lübbecke and Desrosiers,, 2005) to address the computational challenges associated with the large decision variables in the proposed formulation.
+The strength of CG lies in its ability to break down a complex problem into manageable subproblems while iteratively refining the solution.
+The CG algorithm decomposes a large-scale optimization problem into two components: a master problem and a subproblem (see Fig.
+2 (a)).
+The master problem, typically formulated as a linear program (Vanderbeck,, 2005), selects the best combination of columns (also referred to as paths or candidate solutions) from a feasible set $\Omega$.
+Since $\Omega$ is generally too large to consider in full, a restricted master problem is solved over a smaller subset $\Omega^{\prime}\subset\Omega$.
+Solving this restricted problem yields a tentative solution along with dual variables.
+In a linear program, each constraint in the primal problem corresponds to a variable in the dual problem, known as a dual variable (Bachem et al.,, 1992).
+Intuitively, a dual variable represents the rate at which the optimal value of the primal objective would improve if the associated constraint were relaxed by one unit.
+The subproblem uses these dual variables to identify new columns that could improve the current solution.
+Guided by the duals, the subproblem searches $\Omega$ for high-value candidates missing from $\Omega^{\prime}$.
+These candidates are added to $\Omega^{\prime}$, and the master problem is re-solved.
+For example, the dual variables associated with the restricted master problem constraints in Eqs.
+(18)–(19) are used in the objective function of the subproblem defined in Eq.
+(22).
+This iterative process continues until no further columns can be found that would improve the objective value of the master problem.
+The overall performance of the CG algorithm heavily depends on how effectively the subproblem is solved, particularly as the number of agents and tasks increases.
+In such cases, identifying promising solution candidates to expand the restricted set $\Omega^{\prime}$ becomes computationally challenging.
+To address this limitation, we propose a clustering-based heuristic method, coined Clustering-based Column Generation (CCG), to solve the subproblem method efficiently.
+CCG tackles the subproblem one agent at a time, focusing on a small, representative subset of tasks, denoted by $\mathcal{T}^{\prime}$ (see Fig.
+2 (b)).
+This subset is constructed by first clustering the original task set $\mathcal{T}$ based on task time-windows, and then uniformly sampling from these clusters to form $\mathcal{T}^{\prime}$.
+This targeted reduction in task space allows for more scalable and efficient subproblem solving.
+The detailed methodology behind CCG is presented in the following section.
+Master ProblemDual ValuesSubproblemSubset SolutionPool ($\Omega^{\prime}\subset\Omega$)Solution Pool ($\Omega$)
+(a) Block diagram of Column Generation (CG)
+Master ProblemSubset SolutionPool ($\Omega^{\prime}\subset\Omega$)Solution Pool ($\Omega$)Dual ValuesAvailable TasksCreate ClustersSingle-agentMIPSampled task subset($\mathcal{T^{\prime}}$)SubproblemTask Set ($\mathcal{T}$)
+(b) Block diagram of the proposed Clustering Column Generation (CCG)
+Figure 2: (a) Column Generation (CG) decomposes an optimization problem into a master problem and a subproblem.
+The master problem selects the best combination of solutions from a solution pool ($\Omega$), while the subproblem uses dual values from the master to generate promising new candidates.
+Since $\Omega$ can be very large the CG operates on a smaller subset of solutions, i.e.
+$\Omega\subset\Omega^{\prime}$.
+This process iterates until no improving solutions are found or a termination condition is met.
+(b) We propose Clustering-based Column Generation (CCG), which incorporates a clustering heuristic to efficiently solve the subproblem.
+To reduce computational cost, tasks in the orignal task set $\mathcal{T}$, are first grouped into clusters using $k$-means based on time windows.
+For each agent, a MIP with a time limit is solved on a reduced subset of tasks $\mathcal{T^{\prime}}$, obtained by uniformly sampling from these clusters.
+The resulting solutions are added to the subset solution pool ($\Omega^{\prime}$) for use in the master problem.
 III-B1 Restricted Master Problem
-This problem aims to identify the best set of solutions from the existing solution pool, resembling a set covering problem.
-Therefore, to solve the restricted master problem, we adopt a set covering problem formulation [27].
-A solution for the set covering problem selects a subset of sets (in our case, task sequence and modes for each agent) from a collection (the solution pool $\Omega^{\prime}$) such that most of the elements (tasks) are covered.
-This must be done while adhering to constraints, such as each task being assigned to only one agent, and maximizing the given objective.
+This problem aims to identify the best set of solutions from the existing solution pool, resembling a set covering problem (Caprara et al.,, 2000).
+Therefore, to solve the restricted master problem, we adopt a set covering problem formulation (Desrosiers and Lübbecke,, 2005).
+A solution for the set covering problem selects a subset of sets (in our case, task sequences and modes for each agent) from a collection (the solution pool $\Omega^{\prime}$) such that most of the elements (tasks) are covered.
+This must be done while adhering to constraints, such as each task being assigned to only one agent and maximizing the given objective.
 The set covering formulation-based restricted master problem finds optimal solutions from a subset of feasible candidate solutions $\Omega^{\prime}$.
 Each candidate in $\Omega^{\prime}$ is a feasible solution that specifies a sequence of addressing different tasks in the associated modes and arrival times.
-The pool $\Omega^{\prime}$ can be initialized by some trivial solutions and can be updated iteratively with better solutions obtained by solving the sub-problem.
+The pool $\Omega^{\prime}$ can be initialized by some trivial solutions and can be updated iteratively with better solutions obtained by solving the subproblem.
 The set covering formulation for our restricted master problem is given by:
@@
 subject to:
-$\begin{split}\sum_{k\in\Omega^{\prime}}a^{k}_{i}\cdot\theta^{k}\leq 1\,\,\forall i\in\mathcal{T}/\{0\},\\ \end{split}$ (18)
+[…]
 $\sum_{k\in\Omega^{\prime}}b^{k}_{i,m}\cdot\theta^{k}\leq 1\,\,\forall i\in\mathcal{T}/\{0\},\forall m\in\mathcal{M}_{i}$ (19)
@@
 The objective function in Eq.
-(17) selects the candidates from $\Omega^{\prime}$ that maximizes the original objective function Eq.
+(17) selects the candidates from $\Omega^{\prime}$ that maximize the original objective function Eq.
 (23).
@@
 (20) ensures the number of candidates selected from $\Omega^{\prime}$ does not exceed the number of available agents.
-We apply a linear relaxation to the integer program defined in Eq.
-(17)-(21) by relaxing the constraint in Eq.
+Typically, a linear relaxation of the original restricted master problem is solved in column generation (Vanderbeck,, 2005).
+This is because the dual of a mixed integer programming (MIP) problem is not well defined, which poses a challenge since dual variables are essential for solving the subproblem.
+In contrast, dual variables are well defined for linear programs (Guzelsoy and Ralphs,, 2007).
+The linear relaxation of the integer program defined in Eqs.
+(17)–(21) is obtained by relaxing the integrality constraint in Eq.
 (21) and allowing $\theta^{k}\in[0,1]$.
-This relaxation enables faster computation, as algorithms like the simplex method [29] can efficiently solve large-scale linear programs in polynomial time.
-This is especially useful with a large number of columns in $\Omega^{\prime}$.
-III-B2 Clustering-based Sub-problem
-In CG, the sub-problem is solved iteratively to populate $\Omega^{\prime}$.
-Solving the sub-problem efficiently with high-quality candidate solutions enhances the CG’s exploration of the solution space and accelerates convergence to better solutions.
+This relaxation also enables faster computation, as algorithms such as the simplex method (Nabli,, 2009) can efficiently solve large-scale linear programs.
+This is particularly advantageous when the restricted column set $\Omega^{\prime}$ contains a large number of variables.
+III-B2 Clustering-based subproblem
+In CG, the subproblem is solved iteratively to populate $\Omega^{\prime}$.
+Solving the subproblem efficiently with high-quality candidate solutions enhances the CG’s exploration of the solution space and accelerates convergence to better solutions.
 However, as the number of tasks and agents increases, the expanded solution space slows down the search for promising candidates.
-To address this, we introduce a clustering method inspired by cluster-first, route-second heuristics commonly used in routing and scheduling [30].
-Our sub-problem solves the MIP problem in Eq.
+To address this, we introduce a clustering method inspired by cluster-first, route-second heuristics commonly used in routing and scheduling (Liu et al.,, 2023; Elango et al.,, 2011).
+Our subproblem solves the MIP problem in Eq.
 (22) for a single agent at a time, on a restricted set of tasks $\mathcal{T}^{\prime}$.
 All tasks in $\mathcal{T}$ are clustered based on time windows.
-We use a K-Means clustering [31] approach to create $n_{c}$ clusters, with $n_{c}$ being a user-defined parameter.
+We use a $k$-means clustering (Bishop,, 2006) approach to create $n_{c}$ clusters, with $n_{c}$ being a user-defined parameter.
 Then $\mathcal{T^{\prime}}$ is generated by uniformly sampling each cluster.
 These modifications allow CCG to converge to higher-quality solutions faster than the baseline, as demonstrated in Section IV.
-The sub-problem creates a solution for a single agent setup with a restricted task pool $\mathcal{T}^{\prime}$. The formulation for the sub-problem is:
-$\begin{split}\underset{x_{i\scriptsize{\xrightarrow{}}j},y_{i,m}}{\max}\,\sum_{r=0}^{R}\sum_{i\in\mathcal{T}^{\prime}}\sum_{j\in\mathcal{T}^{\prime}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x_{i\scriptsize{\xrightarrow{}}j},y_{i,m},s_{i})-\\ \sum_{i\in\mathcal{T}^{\prime}/\{0\}}(\gamma_{i}\cdot\sum_{j\in\mathcal{T}^{\prime}}x_{i\scriptsize{\xrightarrow{}}j})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}\sum_{m\in\mathcal{M}_{i}}(\beta_{i}\cdot y_{i,m}),\end{split}$ (22)
+The subproblem creates a solution for a single-agent setup with a restricted task pool $\mathcal{T}^{\prime}$. The subproblem formulation is:
+$\begin{split}\underset{x_{i\scriptsize{\xrightarrow{}}j},y_{i,m}}{\max}\,\sum_{r=0}^{R}\sum_{i\in\mathcal{T}^{\prime}}\sum_{j\in\mathcal{T}^{\prime}}\sum_{m\in\mathcal{M}_{i}}\lambda_{r}\cdot f_{r}(x_{i\scriptsize{\xrightarrow{}}j},y_{i,m},s_{i})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}(\gamma_{i}\cdot\sum_{j\in\mathcal{T}^{\prime}}x_{i\scriptsize{\xrightarrow{}}j})-\sum_{i\in\mathcal{T}^{\prime}/\{0\}}\sum_{m\in\mathcal{M}_{i}}(\beta_{i}\cdot y_{i,m}),\end{split}$ (22)
 subject to constraints Eq.
 (2)-(16) and the number of agents is restricted to one.
-The objective function of the sub-problem, Eq.
+The objective function of the subproblem, Eq.
 (22), is a modified variant of Eq.
 (1), tailored for a single-agent setup and augmented with dual variables from the restricted master problem.
-Here, $\gamma_{i}$ and $\beta_{i,m}$ are non-negative dual of the restricted master problem corresponding to constraints Eq.
+Here, $\gamma_{i}$ and $\beta_{i,m}$ are non-negative duals of the restricted master problem corresponding to constraints Eq.
 (18) and Eq.
@@
 These dual variables are used to identify promising new candidate solutions.
-It should be noted that sub-problem The solution of the sub-problem is added to $\Omega^{\prime}$.
-Algorithm 1 The clustering-based column generation algorithm
-[…]
+Dual variables can be directly obtained from the LP relaxation solution provided by the optimizer (Laborie et al.,, 2018).
+It should be noted that the solution of the subproblem is added to $\Omega^{\prime}$.
+Algorithm 1 Clustering-based Column Generation (CCG)
+1:  Input: Problem data ($\mathcal{A},\mathcal{T},\mathcal{M},T_{H},Q$ and associated data for nodes), user preference $\lambda$, and algorithm parameters ($n_{c},|\mathcal{T^{\prime}}|,\text{max-iterations},\text{time-limit}$).
 2:  Output: Selected solution for each agent from $\Omega^{\prime}$
@@
 14:    $\mathcal{T^{\prime}}\leftarrow$ Uniformly sample tasks from $\mathcal{T}_{c}$
-15:        $\mathcal{S}\leftarrow$ Solve the sub-problem using Eq. (22)
+15:    $\mathcal{S}\leftarrow$ Solve the subproblem using Eq. (22)
 16:    if $\mathcal{S}$ is not empty then
@@
 25:  if Terminate then
-26:     Solve Eq (17)-(20) subject to $\theta^{k}=\{0,1\}$
+26:   Solve Eq. (17)-(20) subject to $\theta^{k}=\{0,1\}$
 27:  end if
@@
 (2)-(16) are satisfied.
-From line 5, the restricted master problem and sub-problem are solved iteratively to generate new solutions.
-The restricted master problem is updated in each iteration based on $\Omega^{\prime}$ (line 6), and the sub-problem’s objective is guided by the duals obtained from the restricted master.
+From line 5, the restricted master problem and subproblem are solved iteratively to generate new solutions.
+The restricted master problem is updated in each iteration based on $\Omega^{\prime}$ (line 6), and the subproblem’s objective is guided by the value of dual variables obtained from the restricted master.
 A list $\mathcal{T}_{v}$ is initialized to track visited tasks (line 7).
-The sub-problem is solved iteratively for each agent in $\mathcal{A}$ (lines 8-20).
+The subproblem is solved iteratively for each agent in $\mathcal{A}$ (lines 8-20).
 In the first iteration, task clusters $\mathcal{T}_{c}$ are initialized to $\mathcal{T}_{c}^{init}$ (line 10).
 For subsequent iterations, clusters are recreated from the remaining tasks (line 12).
-The restricted task set $\mathcal{T^{\prime}}$ is formed by uniformly sampling clusters from $\mathcal{T}_{c}$ (line 14), after which the sub-problem is solved (line 15).
+The restricted task set $\mathcal{T^{\prime}}$ is formed by uniformly sampling clusters from $\mathcal{T}_{c}$ (line 14), after which the subproblem is solved (line 15).
 Feasible solutions with positive reduced costs are added to $\Omega^{\prime}$ (line 17), and $\mathcal{T}_{v}$ is updated with tasks corresponding to the best solutions in $\mathcal{S}$ (line 18).
-This process continues until the stopping criteria—maximum iterations, time limit, or lack of solutions with positive reduced costs—are met (lines LABEL:lst:ln21-23).
+This process continues until the stopping criteria—maximum iterations, time limit, or lack of solutions with positive reduced costs—are met (lines 21-23).
 Upon termination, the restricted master problem is solved to optimality with integrality constraints to obtain the final solution (lines 25-28).
-A time limit is applied to both the restricted master and sub-problems to maintain time efficiency throughout the process.
+A time limit is applied to the restricted master and subproblems to maintain time efficiency during the process.
+The worst-case time complexity of CCG is determined by its components.
+The restricted master problem is a linear program, which, when solved using the simplex method, has a worst-case exponential complexity of $O(2^{m})$ (Voulgaropoulou et al.,, 2019), where $m$ is the number of constraints.
+However, it typically runs in polynomial time in practice (Voulgaropoulou et al.,, 2019).
+The subproblem, solved as an MIP using a branch-and-bound solver, also possesses exponential complexity (Huang et al.,, 2021).
+Additionally, the $k$-means clustering step has a worst-case complexity of $O(n^{d})$, where $n$ is the number of nodes and $d$ is the data dimension (Arthur and Vassilvitskii,, 2006).
+However, in practical implementations, $k$-means typically converges in $O(nkT)$ time (Arthur and Vassilvitskii,, 2006), where $k$ is the number of clusters and $T$ is the number of iterations.
+With a fixed iteration limit, its runtime becomes effectively $O(1)$ per instance.
+Since both the master and subproblem have fixed time limits, their complexity per iteration is effectively $O(1)$.
+With an efficient implementation of $k$-means that also runs in bounded time, the overall time complexity is reduced to $O(I)$, where $I$ is the number of iterations until the convergence or time-out of the CCG.
+The space complexity depends on the size of the solution pool, leading to $O(|\Omega^{\prime}|)$.
 III-C M3RS for Multi-Robot Disinfection
 We consider a fleet of surface disinfection robots deployed in hospital-like environments to provide around-the-clock disinfection.
-Specifically, we use G-robots [32], which are far-UV-based mobile manipulators designed for surface disinfection.
+Specifically, we use G-robots (Mehta et al.,, 2022), which are mobile manipulators equipped with far-UVC emitters and are designed for surface disinfection (see Fig.
+7a).
 A disinfection mission is a collection of disinfection tasks spread in an environment such as a hospital.
@@
 Each task can be disinfected in a particular set of modes, where the modes have different disinfection dosage options for that task.
-The disinfection dosage for UV-based surface disinfection is product of UV exposure duration and irradiance of the UV light source [1].
-A higher disinfection dosage would require a longer exposure duration.
-Typically, the disinfection dosage is expressed in terms of percentage of population reduction.
-For example, a dosage of $D_{90}$ implies $90\%$ reduction of pathogenic population.
-A higher reduction percentage requires stronger disinfection dose which in turn leads to longer exposure times.
-Longer exposure times further implies higher energy consumption.
+The disinfection dosage ($D$ in $\frac{\text{Watt}\cdot s}{m^{2}}$) for UV-based surface disinfection is the product of UV exposure duration ($E_{t}$ in $s$) and irradiance of the UV light source ($I_{r}$ in $\frac{\text{Watt}}{m^{2}}$), given by $D=E_{t}\cdot I_{r}$ (Mehta et al.,, 2023).
+A higher disinfection dosage ($D$), requires a longer exposure time ($E_{t}$), given the fixed irradiance of the UV light source ($I_{r}$).
+Disinfection dosage is typically expressed as the percentage reduction in pathogen populations and is pre-calculated for various pathogen types (Mehta et al.,, 2023).
+For example, a dosage of $D_{90}$ implies a $90\%$ reduction of the pathogenic population.
+A higher reduction percentage requires a stronger disinfection dose, which in turn leads to longer exposure times.
+Longer exposure times further imply higher energy consumption.
 We consider four disinfection modes that are $D_{99.9999}$, $D_{99.99}$, $D_{99}$, and $D_{90}$.
-Here, $D_{99.9999}$ is the strongest disinfection dose with the highest disinfection time and energy consumption as compared to $D_{90}$ [1].
+Here, $D_{99.9999}$ is the strongest disinfection dose with the highest disinfection time and energy consumption as compared to $D_{90}$ (Mehta et al.,, 2023).
 The set of modes can be determined by the acceptable disinfection levels for a surface.
-So, each task has an associated set of acceptable disinfect doses or modes.
+So, each task has an associated set of acceptable disinfectant doses or modes.
 The assignment of robots to different tasks and the corresponding modes will be determined by solving M3RS.
-For disinfection applications, we have two objectives: i) the number of tasks completed and ii) overall disinfection quality.
+The multi-objective formulation and execution mode decision variables in M3RS enable trade-offs between disinfection quality and number of tasks completed, based on mission constraints and user preferences.
+For disinfection applications, we have two objectives: i) the number of tasks completed and, i.e.
+$f_{1}$, ii) overall disinfection quality i.e.
+$f_{2}$.
 Higher disinfection quality requires more time and resources, limiting availability for other tasks.
@@
 Our formulation creates schedules by using user-specified trade-off requirements.
-The formulation of M3RS for disinfection application is:
-$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m}}{\max}\,\frac{1}{|\mathcal{T}|}\cdot\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}(\lambda)\cdot x^{k}_{i{\xrightarrow{}}j}+(1-\lambda)\cdot\frac{p_{i,m}}{p^{max}}\cdot y^{k}_{i,m,}$ (23)
-subject to constrains in Eq.
-(3)-(16).
-Here, $p_{i,m}$ indicates the quality of disinfection when completing task $i$ in mode $m$, $p^{max}$ is the highest possible quality of disinfection, and $\lambda$ represents the user-specified preference.
-When $\lambda=1$, the objective ignores quality and focuses on maximizing the number of tasks completed, whereas when $\lambda=0$, the objective exclusively prioritizes maximizing the quality of disinfection.
-This mixed integer optimization model can be solved using a standard branch and cut algorithm.
-However, for larger problem instances, the branch and cut solution has an exploding computation time.
-The CCG algorithm described in Alg.
-1 can be used to effectively compute solutions in such cases.
-(a) A sample layout for synthetic experiment.
-(b) Comparison of idling time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
-(c) Comparison of working time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
-(d) Impact on $SR$ and $DQ$ for different types of time windows. The tasks with time-windows could be critical or non-critical.
-Figure 2: (a) A sample layout of a disinfection mission with different task types.
-(b) & (c) The solution of M3RS make effective use of agents as indicated by lower idling times and higher working times.
-(d) It can be seen that tasks available during the entire mission have a higher completion rate ($SR$) and disinfection quality as well.
+The overall objective of M3RS–composed of $f_{r}(\cdot),\penalty 10000\ r=1,2$, in Eq.
+(1)–for the disinfection application is given by:
+$\underset{x^{k}_{i{\xrightarrow{}}j},y^{k}_{i,m}}{\max}\,\frac{1}{|\mathcal{T}|}\cdot\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}(\lambda)\cdot f_{1}+(1-\lambda)\cdot f_{2},$ (23)
+subject to the constraints in Eqs. (2)–(16). This $f_{1}(\cdot)$ and $f_{2}(\cdot)$ are defined as follows:
+$f_{1}=x^{k}_{i\rightarrow j},$ (24)
+which maximizes the quantity of tasks completed, and
+$f_{2}=\frac{p_{i,m}}{p^{\max}}\cdot y^{k}_{i,m},$ (25)
+which maximizes the quality of completed tasks.
+Here, $p_{i,m}$ denotes the reward representing the disinfection quality for task $i$ executed in mode $m$ (defined in Sec.
+IV-B2), $p^{\text{max}}$ is the maximum achievable disinfection quality reward, and $\lambda$ captures the user-defined preference.
+See Table I for a summary of the notation.
+When $\lambda=1$, the objective prioritizes maximizing the number of tasks completed, disregarding disinfection quality.
+Conversely, when $\lambda=0$, it exclusively focuses on maximizing disinfection quality.
+Users can adjust $\lambda$ based on mission priorities.
+Lower values suit missions with critical areas like ICUs or patient care rooms.
+Higher values are appropriate for missions with less critical spaces, e.g., waiting rooms.
+This mixed integer optimization model can be solved using a standard branch-and-cut algorithm.
+However, for larger problem instances, the branch-and-cut has an exponentially increasing computation time.
+The CCG algorithm, Alg.
+1, can be used in such cases.
 IV Evaluation
-In this section, we evaluate the the M3RS problem and propesed solution strategy of CCG.
+In this section, the M3RS problem and the proposed solution of CCG are evaluated.
+The evaluation is conducted for multi-robot disinfection scenarios.
 First, the metrics used for evaluation are defined.
-Second, the impact of using task-level quality as a decision variable is studied by comparing M3RS with the classic routing and scheduling in fixed (RS-F) disinfection mode on synthetically generated data.
-Then, the proposed CCG scheme is compared with a mixed integer programming (MIP) solver.
+Then, the impact of using task-level disinfection quality as a decision variable is analyzed.
+This is done by comparing M3RS with a baseline that performs classic routing and scheduling in fixed (RS-F) disinfection modes.
+Fixed modes are used since existing MRTA methods do not treat quality as a decision variable, and no relevant baselines exist.
+The comparison is performed on synthetically generated multi-robot disinfection missions.
+Additioannly, the proposed CCG scheme is compared with a mixed integer programming (MIP) solver on synthetically generated missions.
 Finally, hardware and simulated case studies are presented.
-In each evaluation, multiple problem instances are generated. Each problem instance is characterized by mission time which is given by:
-$T_{H}=\frac{t_{avg}\times\text{\# Tasks}}{\eta\times\text{Fleet size}}$ (24)
-Here, $\eta$ is the load factor that is used to simulate missions with different $T_{H}$.
-A higher $\eta$ leads to tighter mission times.
+In each evaluation, multiple problem instances are generated. Each instance is characterized by mission time, given by:
+$T_{H}=\frac{t_{avg}\times\text{\# Tasks}}{\eta\times\text{Fleet size}}.$ (26)
+Here, $T_{H}$ represents the mission time, i.e., the total time available to complete all tasks, while $\eta$ is the load factor used to simulate missions with varying $T_{H}$.
+The load factor is commonly used in MRTA to represent mission workload relative to the robot’s available capacity (Botros et al.,, 2023).
+We use the load factor to model varying mission durations, where a higher $\eta$ corresponds to tighter time constraints for the mission.
 The average task completion time is given by $t_{avg}$.
 In our experiments, $t_{avg}$ is considered to be 4.5 minutes.
-We assume that the robot travels at a fixed speed of $0.5\,{m}/{s}$.
-We use the MIP solver from IBM’s CPLEX [33].
-All the evaluations are carried out on machine equipped with an Intel i9 processor with 32 GB RAM.
+We assume that each robot travels at a fixed speed of 0.5m/s0.5\,{m}/{s}.
+We use the MIP solver from IBM’s CPLEX (Laborie et al.,, 2018).
+All evaluations are done on a machine with an Intel i9 processor and 32GB RAM.
 IV-A Evaluation Metrics
-We define some evaluation metrics to compare the performance of M3RS and RS-F (Sec. IV-B). The metrics are:
-$SR\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}}{|\mathcal{T}|}$ (25)
-$DQ\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}p_{i,m}\cdot y^{k}_{i,m}}{|\mathcal{T}|\cdot p^{max}}$ (26)
-$MSI\,=\,\frac{SR+DQ}{2}$ (27)
-$ATQ\,=\,\frac{DQ}{SR}$ (28)
-$QBR\,=\,\frac{DQ-DQ^{*}}{(\bar{Q}_{exp}-\bar{Q}_{exp}^{*})+(\bar{M}_{exp}-\bar{M}_{exp}^{*})}$ (29)
+We define six evaluation metrics to compare the performance of M3RS against other methods. The metrics are:
+$SR\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{j\in\mathcal{T}}x^{k}_{i{\xrightarrow{}}j}}{|\mathcal{T}|},$ (27)
+$DQ\,=\,\frac{\sum_{k\in\mathcal{A}}\sum_{i\in\mathcal{T}}\sum_{m\in\mathcal{M}_{i}}p_{i,m}\cdot y^{k}_{i,m}}{|\mathcal{T}|\cdot p^{max}},$ (28)
+$MSI\,=\,\frac{SR+DQ}{2},$ (29)
+$ATQ\,=\,\frac{DQ}{SR},$ (30)
+$QBR\,=\,\frac{DQ-DQ^{*}}{(\bar{Q}_{exp}-\bar{Q}_{exp}^{*})+(\bar{M}_{exp}-\bar{M}_{exp}^{*})}.$ (31)
 The success rate ($SR$) measures the number of tasks completed successfully (Eq.
@@
 (30)).
-The quality benefit ratio ($QBR$) compares a target solution with a baseline solution in terms of the gains made in disinfection quality relative to the resource usage (Eq.
+The quality-benefit ratio ($QBR$) compares a target solution with a baseline solution in terms of the gains made in disinfection quality relative to the resource usage (Eq.
 (31)).
@@
 For the target solution, $\bar{Q}_{exp}$ and $\bar{M}_{exp}$ represent the fleet’s average energy expenditure and makespan.
-The RS-F solution serves as the baseline with $DQ^{*}$ as disinfection quality, $\bar{Q}_{exp}^{*}$ as average energy, and $\bar{M}_{exp}^{*}$ as makespan.
-In Sec IV-C, we use relative optimal gap ($OG$) and the objective values (Eq.
-(23)) are also as a metric.
-The $OG$ measures the difference between the best known bound on the objective value ($o^{*}$) and the current solution’s objective value $o$, is also reported.
-MIP solvers typically provide this bound by solving a relaxed version of the problem and is given by:
-$OG\,=\,\frac{o^{*}-o}{o^{*}}$ (30)
-IV-B M3RS vs RS-F
+The RS-F solution serves as the baseline with $DQ^{*}$ as disinfection quality, $\bar{Q}_{exp}^{*}$ as average energy, and $\bar{M}_{exp}^{*}$ as makespan, where makespan is time taken by robot to complete its assigned task during the mission.
+The objective values (Eq.
+(23)) are also used as a metric.
+This is achieved by measuring the difference between the best-known bound on the objective value ($o^{*}$) and the current solution’s objective value, $o$.
+MIP solvers typically provide the bound by solving a relaxed version of the problem.
+This metric is presented as the relative optimality gap ($OG$) (Kaliszewski and Miroforidis,, 2021):
+$OG\,=\,\frac{o^{*}-o}{o^{*}}\cdot 100\,\%.$ (32)
+Quantity Value
+Layout area $150,200,225\,m^{2}$
+Number of tasks 306030-60
+Proportion of critical tasks $15-25\%$
+Avg. task completion time 4.5mins4.5\,mins
+Task surface area range $0.5-2.2\,m^{2}$
+$\eta$ 1
+Fleet size 4
+$Q$ $6\text{Ah}$
+Table II: A list of parameters used to simulate disinfection missions to assess the impact of quality. A mission consists of disinfection tasks spread across a space, see Fig. 3.
+Figure 3: A sample physical layout of locations of different disinfection tasks. Here, the color coding is used to present different categories of tasks.
+IV-B M3RS vs RS-F (Routing and Scheduling with Fixed modes)
+In this experiment, we assess the effectiveness of our proposed M3RS formulation in multi-robot disinfection scenarios.
+The results demonstrate that M3RS outperforms baseline approaches across key performance metrics.
 IV-B1 Baselines
-We assess the impact of the M3RS formulation, which incorporates quality levels as a decision variable through modes, compared to RS-F which is routing and scheduling with fixed execution modes.
-The RS-F formulation follows the vehicle routing problem with time windows [34], and we consider two variants: i) RS-F (Min), which uses the mode with the lowest disinfection quality, and ii) RS-F (Max), which uses the mode with the highest quality.
-For M3RS we use different values of user prefrence i.e.
+We evaluate the impact of the M3RS formulation, which integrates quality levels as a decision variable through multiple execution modes, against RS-F, a routing and scheduling approach with fixed execution modes.
+RS-F follows the vehicle routing problem with time windows (Kallehauge et al.,, 2005) and does not account for task-level quality.
+We consider two RS-F variants: i) RS-F (Min), which defaults to the lowest disinfection dosage, and ii) RS-F (Max), which defaults to the highest disinfection dosage.
+For M3RS we use different values of user preference i.e.
 $\lambda$.
@@
 Both M3RS and RS-F (Min and Max) are solved using an MIP solver with a 10-minute time limit per problem instance.
-The evaluation metrics include $SR$, $DQ$, $MSI$ $ATQ$, and $QBR$, see Eq (25)-(29).
+The evaluation metrics include $SR$, DQDQ, MSIMSI ATQATQ, and $QBR$, see Eq.
+(27)-(31).
 RS-F (Min) is used as the baseline for $QBR$ calculation since it operates in the lowest disinfection mode, allowing us to measure the improvement in disinfection quality relative to resource expenditure.
@@
 IV-B2 Synthetic Data Generation
+The synthetic data simulates realistic hospital missions with rectangular layouts of varying sizes.
+For each layout type, $10$ random problem instances are generated and solved.
+A sample layout is shown in Fig.
+3.
+We consider four disinfection modes, indexed by $m$, where a lower $m$ indicates higher disinfection quality.
+Specifically, $m=0$ corresponds to $D_{99.9999}$ (highest quality), followed by $D_{99.99}$ ($m=1$), $D_{99}$ ($m=2$), and $D_{90}$ ($m=3$, lowest quality).
+For task $i$, the mode set ($\mathcal{M}_{i}$) is randomly assigned between one and four disinfection modes.
+The associated quality reward (see Eq.
+(23)) for each mode is defined as $p_{i,m}=2^{-m}$ for $m\in\{0,1,2,3\}$.
+For example, selecting the highest quality mode ($m=0$) for task $i$ yields the maximum reward of $1$, while the lowest quality mode ($m=3$) yields a reward of $0.125$.
+For each disinfection task, the target surface is assumed to be planar, with an area randomly assigned between $0.5$ and $2.2\,m^{2}$.
+The required dosage, exposure time, and energy consumption for disinfecting the surface are computed based on UV disinfection dosage and energy models (Mehta et al.,, 2022).
+A summary of the experimental setup is provided in Table II.
+Each problem instance involves a fleet of four robots.
+We set the value of load factor $\eta=1$ for this experiment.
+This ensures that for the disinfection mission there is sufficient time to address the tasks.
+The mission time ($T_{H}$) is calculated using Eq.
+(26).
+Tasks are categorized as critical or non-critical.
+In each problem instance, a subset of critical tasks, representing $15-25\%$ of the total tasks, must be completed within a specified time window using the highest disinfection dosage, $D_{99.9999}$.
+These are crucial surfaces, such as those in patient rooms and ICUs.
+The inequality in Eq.
+(2) is replaced by an equality constraint for critical tasks.
+Non-critical tasks, like those in waiting areas, washrooms, and office spaces, can be skipped but contribute to overall hygiene if addressed.
+These tasks may have fixed time windows, and deadlines, or be available throughout the mission.
+All time windows are uniformly sampled from the mission time horizon.
 […]
-Table II: Experiment parameters for testing impact of quality.
-Quantity Value
-Layout area $150,200,225\,m^{2}$
-Number of tasks $30-60$
-Proportion of Critical tasks $15-25\%$
-Avg. task completion time $4.5\,mins$
-Task surface area range $0.5-2.2\,m^{2}$
-$\eta$ 1
-Fleet size 4
-[…]
-Table III: Comparison of performance metrics for synthetic problem instances.
-The colors indicate the first, second, and third ranking for a given metric.
-The solutions of M3RS perform well on both quality and quantity objectives as indicated by different metrics.
-The number of tasks range from $30-60$ and the robot fleet size is four.
-Metric RS-F (Min) RS-F (Max) M3RS-0.01 M3RS-0.1 M3RS-0.5 M3RS-0.9
-$SR$ $1.0\pm 0.0$ $0.847\pm 0.082$ $0.906\pm 0.078$ $0.912\pm 0.072$ $0.941\pm 0.068$ $0.973\pm 0.042$
-$DQ$ $0.452\pm 0.031$ $0.757\pm 0.083$ $0.820\pm 0.078$ $0.829\pm 0.097$ $0.789\pm 0.095$ $0.750\pm 0.124$
-$MSI$ $0.726\pm 0.016$ $0.802\pm 0.082$ $0.863\pm 0.072$ $0.870\pm 0.079$ $0.865\pm 0.073$ $0.862\pm 0.08$
-$ATQ$ $0.452\pm 0.031$ $0.894\pm 0.024$ $0.906\pm 0.063$ $0.909\pm 0.070$ $0.839\pm 0.082$ $0.768\pm 0.104$
-$QBR$ N/A $0.331\pm 0.064$ $0.364\pm 0.082$ $0.391\pm 0.117$ $0.331\pm 0.104$ $0.297\pm 0.129$
-Table IV: Comparison of $OG$ values between MIP solver and CG solutions.
-Colors represent the first, second, and third rankings for the $OG$ metric.
-CCG solutions achieve better $OG$ values as compared to the MIP solver for the same computation time limit.
-Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.9$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.9$
-120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
-MIP $0.131\pm 0.034$ $0.278\pm 0.113$ $0.247\pm 0.046$ $0.353\pm 0.024$ $0.253\pm 0.032$ $0.290\pm 0.061$ $0.570\pm 0.065$ $0.488\pm 0.039$
-CCG ($K=2$) $0.136\pm 0.050$ $0.110\pm 0.024$ $0.218\pm 0.024$ $0.300\pm 0.018$ $0.116\pm 0.015$ $0.177\pm 0.039$ $0.455\pm 0.011$ $0.380\pm 0.006$
-CCG ($K=4$) $0.126\pm 0.009$ $0.121\pm 0.034$ $0.211\pm 0.017$ $0.302\pm 0.023$ $0.097\pm 0.045$ $0.179\pm 0.021$ $0.434\pm 0.013$ $0.363\pm 0.014$
-CCG ($K=6$) $0.110\pm 0.025$ $0.132\pm 0.025$ $0.208\pm 0.010$ $0.305\pm 0.021$ $0.107\pm 0.025$ $0.180\pm 0.021$ $0.453\pm 0.026$ $0.358\pm 0.029$
-CCG ($K=8$) $0.133\pm 0.031$ $0.102\pm 0.022$ $0.220\pm 0.019$ $0.308\pm 0.009$ $0.106\pm 0.047$ $0.184\pm 0.020$ $0.446\pm 0.019$ $0.355\pm 0.025$
-Method Number of Tasks and Robots, $\eta=1.0$, $\lambda=0.1$ Number of Tasks and Robots, $\eta=2.0$, $\lambda=0.1$
-120-4 150-5 175-5 200-6 120-4 150-5 175-5 200-6
-MIP $0.124\pm 0.032$ $0.181\pm 0.021$ $0.153\pm 0.014$ $0.215\pm 0.019$ $0.409\pm 0.031$ $0.586\pm 0.013$ $0.5200\pm 0.026$ $0.602\pm 0.021$
-CCG ($K=2$) $0.103\pm 0.018$ $0.144\pm 0.037$ $0.188\pm 0.022$ $0.205\pm 0.016$ $0.357\pm 0.037$ $0.483\pm 0.015$ $0.425\pm 0.018$ $0.514\pm 0.021$
-CCG ($K=4$) $0.116\pm 0.018$ $0.126\pm 0.019$ $0.156\pm 0.029$ $0.197\pm 0.020$ $0.310\pm 0.023$ $0.491\pm 0.016$ $0.436\pm 0.021$ $0.517\pm 0.018$
-CCG ($K=6$) $0.132\pm 0.027$ $0.150\pm 0.034$ $0.162\pm 0.032$ $0.201\pm 0.020$ $0.349\pm 0.031$ $0.483\pm 0.012$ $0.418\pm 0.009$ $0.520\pm 0.014$
-CCG ($K=8$) $0.116\pm 0.019$ $0.134\pm 0.015$ $0.198\pm 0.010$ $0.181\pm 0.027$ $0.336\pm 0.011$ $0.491\pm 0.012$ $0.424\pm 0.015$ $0.502\pm 0.025$
 IV-B3 Results
-The mean and standard deviation of the performance metrics for solutions of RS-F and M3RS variants are presented in Table III.
-RS-F (Min) achieves the highest success rate ($SR$), addressing all tasks.
-This is because the lowest disinfection dosage is used by default, reducing disinfection time and increasing $SR$, but at the cost of lower disinfection quality ($DQ$).
-Conversely, RS-F (Max) provides higher $DQ$ but lower $SR$.
-The higher disinfection quality requires more time, limiting agents’ ability to complete other tasks.
-The solutions of M3RS outperform the RS-F baselines in the combined metric ($MSI$), offering a better balance between $SR$ and $DQ$.
-The solutions of M3RS admit better $DQ$ and $ATQ$ scores.
-The $QBR$ scores show that solutions of M3RS offer better quality improvement relative to the additional resource expenditure (energy and time).
-This suggests that considering disinfection quality through variable disinfection modes improves overall disinfection quality while maintaining or improving task completion rates.
+Fig.
+4 presents the box-plots of performance metrics for RS-F and M3RS variants.
+RS-F (Min) uses the lowest disinfection mode.
+This minimizes execution time and maximizes success rate ($SR$), due to reduced task completion time.
+However, it results in the lowest disinfection quality ($DQ$).
+RS-F (Max) defaults to the highest disinfection mode.
+This increases execution time and prioritizes disinfection quality.
+However, it lowers the success rate due to longer disinfection times.
+M3RS ($\lambda=0.01$) improves disinfection quality by 10% over RS-F (Max) and 38% over RS-F (Min) while maintaining a strong balance between efficiency and quality.
+In terms of the combined metric ($MSI$), M3RS ($\lambda=0.01$) achieves 10% and 15% gains over RS-F (Max) and RS-F (Min), respectively, demonstrating its ability to optimize both success rate and disinfection quality.
+M3RS solutions also achieve higher Average Task Quality ($ATQ$), with a 3% gain over RS-F (Max) and 46% over RS-F (Min).
+The Quality-Benefit Ratio ($QBR$) improves by 6% compared to RS-F (Max), indicating that M3RS provides better disinfection quality relative to resource expenditure (energy and time).
+These results indicate that incorporating variable disinfection modes yields better overall disinfection quality while maintaining or improving task completion rates.
+To further ascertain the gains made by task execution modes, we conducted a statistical comparison of success rate $SR$ and disinfection quality $DQ$ between RS-F (Max) and M3RS ($\lambda=0.01$) using Welch’s t-test (Sakai and Sakai,, 2018).
+This test was applied to evaluate whether the differences observed in each metric between the two variants are statistically significant.
+For the $SR$, the p-value was $7.74\times 10^{-6}$, and for the $DQ$, the p-value was $5.08\times 10^{-6}$.
+Both results indicate statistically significant differences at the $95\%$ confidence level, demonstrating that M3RS ($\lambda=0.01$) achieves significant improvements over RS-F (Max) in these metrics.
+(a) Box-plot comparing idling time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
+(b) Box-plot comparing working time to $T_{H}$ ratio for M3RS and RS-F for different number of tasks and a fleet size of 4.
+(c) Box-plot depicting impact on SRSR and $DQ$ for different types of time windows. The tasks with time-windows could be critical or non-critical.
+Figure 5: (a) & (b) The solution of M3RS makes effective use of agents as indicated by lower idling times and higher working times.
+(c) Further, tasks available during the entire mission have a higher task completion rate and disinfection quality.
 We further analyze the missions by looking at the ratio of idle time to task servicing time, as well as the $SR$ and $DQ$ for different time-window types (see Fig.
 5 (a)-(c)).
-The solution of M3RS show less idle time and more working time, indicating better use of mission time, see Fig.
+The solution of M3RS shows less idle time and more working time, indicating better use of mission time; see Fig.
 5 (a)-(b).
-Using variable disinfection modes likely saves time and energy, which can be spent on disinfecting more tasks or improving quality for others, depending on the user preference i.e.
+Use of variable disinfection modes promotes better use of time and energy, which could be spent on disinfecting more tasks or improving quality for others, depending on the user preference i.e.
 $\lambda$.
-RS-F (Max) shows higher idle time and lower working time than M3RS, as disinfection in the highest modes requires robots to arrive at or before a task’s time window.
-Early arrivals lead to waiting, limiting task completion and resulting in lower SR and DQ compared to M3RS.
+In most cases, RS-F (Max) shows higher idle time and lower working time than M3RS, as disinfection in the highest modes requires robots to arrive at or before a task’s time window.
+Early arrivals lead to waiting, which limits task completion and results in lower SR and DQ compared to M3RS.
+For missions with more than 50 tasks, we observe that for certain $\lambda$ values, M3RS shows slightly higher average idling times and lower average working times compared to RS-F (Max).
+This could be attributed to the increased problem size with more tasks and may be improved with more efficient solution strategies.
 When examining $SR$ and $DQ$ across time-window types, tasks available throughout the mission have the best performance, while those with limited time windows perform worse, see Fig.
 5 (c).
-This suggests that relaxing time constraints where possible could lead to better overall performance.
+This suggests that relaxing time constraints, where possible, could lead to better overall performance.
+Overall, the synthetic experiments demonstrate that incorporating quality as a decision variable in M3RS enables more effective planning of disinfection missions, balancing the number of tasks completed with the quality of service within given time and resource constraints.
+While this work focuses on disinfection, M3RS is also applicable to other domains, such as cleaning robots, where balancing cleaning efficiency and time is critical (Megalingam et al.,, 2025), and robot patrolling, which involves managing trade-offs between high- and low-priority areas (Katole et al.,, 2023).
 IV-C Performance of CCG
+We evaluate our clustering-based column generation (CCG) method for solving M3RS by comparing it with a standard MIP solver (Laborie et al.,, 2018).
+The results show that CCG achieves comparable or better performance in less time.
 IV-C1 Baselines
-The performance of solving M3RS with CCG is compared with that of directly solving using MIP solver.
+The performance of solving M3RS with CCG is compared with that of directly solving using an MIP solver.
 For CCG, we report results with varying number of clusters i.e $K\in[2,4,6,8]$.
-The performance is evaluated in terms of the relative optimality gap ($OG$) given by Eq (30).
+The performance is evaluated in terms of the relative optimality gap ($OG$) given by Eq.
+(32).
 For CCG, we use the best bound from the MIP solver to compute the $OG$.
 IV-C2 Setup
-For this, four problem types are considered with a number of tasks and robots mentioned in Table.
-IV.
-For each problem type, 5 instances are generated using the procedure outlined in Sec.
+For this, four problem types are considered with a number of tasks and robots mentioned in Fig.
+6.
+For each problem type, 10 instances are randomly generated using the procedure outlined in Sec.
 IV-B.
 User preference for each instance is sampled from $\lambda\in[0.1,0.9]$, with higher $\lambda$ values prioritizing task completion over disinfection quality.
-We use load factors $\eta=1$ and $\eta=2$, where higher $\eta$ simulates missions with shorter $T_{H}$, providing less time for mission completion.
-The mission time ($T_{H}$) is calculated using Eq (24).
-The time limit for MIP solver and CCG is set to 10 minutes.
-We further report the performance with time limit for MIP set to 25 minutes.
+We use load factors $\eta=1$ and $\eta=2$, where a higher $\eta$ simulates missions with shorter $T_{H}$, providing less time for mission completion.
+The mission time ($T_{H}$) is calculated using Eq.
+(26).
+The time limit for the MIP solver and CCG is set to 10 minutes.
+We further report the performance with a time limit for MIP set to 25 minutes.
+(a) $\eta=1.0$ and $\lambda=0.9$
+(b) $\eta=2.0$ and $\lambda=0.9$
+(c) $\eta=1.0$ and $\lambda=0.1$
+(d) $\eta=2.0$ and $\lambda=0.1$
+Figure 6: (a)-(d) Box-plots comparing the optimality gap ($OG$) of M3RS problems solved by: (i) a standard MIP solver (Laborie et al.,, 2018), and (ii) our Clustering-based Column Generation (CCG) approach.
+CCG is evaluated with different numbers of clusters ($K=2,4,6,8$).
+Both methods are given a 10-minute time limit.
+Box-plots are generated from multiple randomly generated problem instances, varying in $\lambda$ and $\eta$.
+Lower $OG$ values indicate better performance.
+CCG consistently achieves a lower OGOG than the MIP solver within the same time budget.
+Table III: Average optimality gap ($OG$) difference between CCG and the MIP solver across various instances, with a 25-minute time limit for MIP and a 10-minute time limit for CCG.
+Positive values indicate better OGOG for MIP, while negative values favor CCG.
+All values are given in $\%$.
+Colors highlight MIP outperforming CCG and CCG outperforming MIP.
+Our CCG method achieves comparable or better OGOG while using 60% less computation time than the MIP solver.
+Metric $\eta$ $\lambda$ # tasks - # robots.
+1204120-4 1505150-5 1755175-5 2006200-6
+$OG_{CCG}-OG_{MIP}$ $1.0$ $0.9$ $4.1\%$ $7.2\%$ $3.2\%$ $6.0\%$
+$0.1$ $5.0\%$ $3.0\%$ $0.9\%$ $0.6\%$
+$2.0$ $0.9$ $0.7\%$ $15.7\%$ $4.2\%$ $15.4\%$
+$0.1$ $0.1\%$ $8.6\%$ $6.2\%$ $9.2\%$
 IV-C3 Discussion
-The mean and standard deviation of the $OG$ metric for MIP and CCG with time limit set to ten minutes is reported in Table IV.
-The change in $OG$ for MIP solver relative to CCG, with time limit for MIP solver set to 25 minutes is reported in Table V.
-The $OG$ achieved by CCG solutions is consistently better than that of MIP solver with the same time limit, see Table IV.
-Further, it is seen with increased time limit the performance of CCG remains comparable or better than MIP solver while being 2.5×2.5$\times$2.5 × faster.
-This indicates that clustering and sampling approach coupled with column generation provides an effective way to find good solution in shorter time as compared to MIP solver.
-Generally, clustering improves $OG$ values, with more clusters benefiting larger instances.
-Table V: Difference of average optimality gap ($OG$) between CCG and the MIP solver across various instances.
-A 25-minute time limit is set for the MIP solver.
-Positive differences indicate better $OG$ for the MIP solver, while negative values favor CCG.
-Colors represent MIP outperforming CCG and CCG outperforming MIP.
-CCG achieves comparable or better $OG$ while being approximately 2.5×2.5$\times$2.5 × faster than the MIP baseline.
-Metric $\eta$ $\lambda$ $120-4$ $150-5$ $175-5$ $200-6$
-$OG_{CCG}-OG_{MIP}$ $1.0$ $0.9$ $0.038$ $-0.084$ $-0.058$ $-0.077$
-$0.1$ $0.022$ $-0.052$ $-0.007$ $-0.026$
-$2.0$ $0.9$ $0.003$ $-0.154$ $-0.074$ $-0.150$
-$0.1$ $-0.029$ $-0.080$ $-0.091$ $-0.105$
+[…]
 IV-D Simulated and Hardware Case Study
-(a) A team of G-robots disinfecting tables.
-(b) The simulated environment with 20 tasks and 2 robots.
-(c) The route sequences for simulated mission with 20 tasks, 2 robots, $T_{H}$ of 0.4 hours and $\lambda=0.9$.
-(d) The route sequence for hardware experiment with $\lambda=0.9$
-(e) The route sequence for hardware experiment with $\lambda=0.1$
-Figure 3: (a) A team of G-robots [32] disinfecting tables.
-(b) The routes of two robots executing various tasks in the simulated environment.
-The route of the first robot is highlighted in white and for the second robot yellow is used.
-(c) The route sequence taken by each robot in the simulated mission.
-For $\lambda=0.9$, maximization of the number of tasks completed has more emphasis than the quality of disinfection.
-As a result, most of the tasks are done in higher modes.
-(d) & (e) The route sequence taken by each robot for the hardware mission with $\lambda=0.9$ and $0.1$.
-For $\lambda=0.9$, the emphasis is on disinfecting all the tables at higher disinfection modes.
-Whereas, for $\lambda=0.1$, the emphasis is on ensuring high-quality disinfection rather than disinfecting all tables.
+In this section, simulation and hardware case studies are presented which apply M3RS in realistic scenarios.
+Videos of the experiments are available on the project page: https://sites.google.com/view/g-robot/m3rs/.
 IV-D1 Simulation and Hardware Experiment Setup
 Multiple disinfection mission instances are generated for a simulated environment.
-Each disinfection mission is characterized by mission time, time windows for each task, and number of robots used in the mission.
+Each disinfection mission is characterized by mission time, time windows for each task, and the number of robots used in the mission.
 The simulated environment consists of $20$ tables as disinfection surfaces spread across two rooms.
-The fleet size considered in the range, $|\mathcal{A}|=2-5$.
+The fleet size considered in the range, $|\mathcal{A}|=2$ to $5$.
 The mission time is calculated using Eq.
-24.
-The entire simulated experiment is conducted in ROS [35] and Gazebo environment [36].
+(26).
+The entire simulated experiment is conducted in ROS (Quigley et al.,, 2009) and the Gazebo environment (Koenig and Hsu,, 2013).
 A simulated mission is depicted in Fig 7 (b).
@@
 7 (a)) for a disinfection mission consisting of six tables.
-Each table can be disinfected in three disinfection modes, i.e.
-$D_{99.99}$, $D_{99}$, and $D_{90}$.
+Each table can be disinfected in three disinfection modes, i.e., $D_{99.99}$, $D_{99}$, and $D_{90}$.
 The entire disinfection mission must be completed in $17$ minutes.
 For both experiments, M3RS is solved for two preferences: $\lambda\,=0.9$ and $0.1$ for each mission.
-The optimization model is solved in CPLEX [33].
-The metrics reported include $SR$, $DQ$, $MSI$, and average lateness of the robot in task completion.
+The optimization model is solved in CPLEX (Laborie et al.,, 2018).
+The reported metrics are $SR$, DQDQ, MSIMSI, and the average lateness of the robot in task completion.
+(a) A team of G-robots disinfecting tables.
+(b) The simulated environment with 20 tasks and 2 robots.
+(c) The route sequence for hardware experiment with $\lambda=0.9$
+(d) The route sequence for hardware experiment with $\lambda=0.1$
+(e) The route sequences for simulated mission with 20 tasks, 2 robots, $T_{H}$ of 0.4 hours and $\lambda=0.9$.
+Figure 7: (a) A team of G-robots (Mehta et al.,, 2022) disinfecting tables.
+(b) The routes of two robots executing various tasks in the simulated environment.
+The route of the first robot is highlighted in white and for the second robot yellow is used.
+(c) & (d) The route sequence taken by each robot for the hardware mission with $\lambda=0.9$ and $0.1$.
+For $\lambda=0.9$, the emphasis is on disinfecting all the tables at higher disinfection modes.
+Whereas, for $\lambda=0.1$, the emphasis is on ensuring high-quality disinfection rather than disinfecting all tables.
+(e) The route sequence taken by each robot in the simulated mission.
+For $\lambda=0.9$, maximization of the number of tasks completed has more emphasis than the quality of disinfection.
+As a result, most of the tasks are done in higher modes.
 IV-D2 Discussion
-The metrics for simulated and hardware case studies are reported in Table VI.
-The example of generated route sequences in simulated experiment is visualized in Fig.
-3 (b)-(c).
+The metrics for simulated and hardware case studies are reported in Table IV.
+The example of generated route sequences in a simulated experiment is visualized in Fig.
+7 (b) & (e), respectively.
 The route sequence for hardware experiments is shown in Fig.
-3 (d)-(e).
-From Table VI, it can be seen that a decent $SR$ is achieved in all missions.
-Even for missions with tight $T_{H}$ the $SR$ is high.
-Therefore, a larger robotic fleet would be necessary to compensate for tighter mission times in order to service as many tasks as possible.
-For missions with shorter $T_{H}$, there is a decline in $DQ$ metric.
+7 (c)-(d).
+From Table IV, it can be seen that a decent $SR$ is achieved in all missions.
+Even for missions with tight $T_{H}$, the $SR$ is high.
+Therefore, a larger robotic fleet would be necessary to compensate for tighter mission times to service as many tasks as possible.
+For missions with shorter $T_{H}$, there is a decline in the $DQ$ metric.
 This is because with tighter mission times, each task has a smaller time window, which means many of the tasks cannot be performed in higher disinfection modes.
 A possible remedy could be to enable robots to collaborate on different tasks.
-Finally, there are some minor delays observed in terms of task completion times for simulation and hardware implementation.
-These delays are caused by slightly longer path navigation times.
-Such delays can be accounted for by introducing chance constraints to model stochasticity in the system.
-Table VI: Results for simulated and hardware experiments.
-Num. of tasks and robots $T_{H}$ (hrs) $\lambda$ $SR$ $DQ$ MSI Avg. Lateness (s)
+Finally, delays in task completion times were observed in both simulation and hardware implementations.
+These are primarily due to longer-than-expected navigation times caused by environmental and planning uncertainties.
+Such delays can be mitigated by incorporating chance constraints to model system stochasticity more effectively.
+Table IV: Results for simulated and hardware experiments.
+# tasks - # robots $T_{H}$ (hrs) $\lambda$ SRSR DQDQ MSI Avg. Lateness (s)
 20-2 0.40 0.9 0.80 0.57 0.63 1.9
@@
 We first propose a mixed-integer linear programming model for M3RS.
-Then, a column generation approach is presented to tackle larger instances with a faster computation time.
 We showcase the utility of M3RS in multi-robot disinfection applications.
-The experiments suggest that our solution for M3RS provides more flexibility to users to select solutions based on desired preferences.
-The current approach is limited as it assumes a deterministic environment, unlike real-world scenarios.
-Generating the entire Pareto front is computationally expensive, particularly for larger instances.
-Our future work will extend M3RS to incorporate elements like stochastic travel times, dynamic arrival of tasks and realistic issues like robot breakdown.
-Further, we aim to tightly integrate the robot path planning with the decision making framework provided by M3RS.
-Finally, we will investigate using deep reinforcement learning techniques to generate good quality solutions in real-time.
-Acknowledgements
+The experiments suggest that our M3RS demonstrates 3%–46% performance improvements over standard task allocation methods across various metrics.
+The solutions for M3RS provide more flexibility to users to select solutions based on desired preferences.
+We also presented a clustering-based column generation approach that solves M3RS instances in $60\%$ lower compute time while maintaining competitive solutions.
+The current approach assumes a deterministic environment and simplified robot energy models.
+Our future work will extend M3RS to incorporate elements like stochastic travel times, dynamic arrival of tasks, realistic robot energy models, and realistic issues like robot breakdown.
+Further, we aim to tightly integrate the robot path planning with the decision-making framework provided by M3RS.
+Finally, we will investigate using deep reinforcement learning techniques to generate good-quality solutions in real-time.
+Acknowledgments
 We would like to thank Mr. Matthew Lisondra for all the support with the experiments.